### v1

# RadMamba: Efficient Human Activity Recognition through Radar-based Micro-Doppler-Oriented Mamba State-Space Model

Abstract
Radar-based Human Activity Recognition (HAR) has emerged as a promising alternative to conventional monitoring approaches, such as wearable devices and camera-based systems, due to its unique privacy preservation and robustness advantages. However, existing solutions based on convolutional and recurrent neural networks, although effective, are computationally demanding during deployment. This limits their applicability in scenarios with constrained resources or those requiring multiple sensors. Advanced architectures, such as Vision Transformer (ViT) and State Space Model (SSM) architectures, offer improved modeling capabilities and have made efforts toward lightweight designs. However, their computational complexity remains relatively high. To leverage the strengths of transformer architectures while simultaneously enhancing accuracy and reducing computational complexity, this paper introduces RadMamba, a parameter-efficient, radar micro-Doppler-oriented Mamba SSM specifically tailored for radar-based HAR. Across three diverse datasets, RadMamba matches the top-performing previous model’s 99.8% classification accuracy on Dataset DIAT with only 1/400 of its parameters and equals the leading models’ 92.0% accuracy on Dataset CI4R with merely 1/10 of their parameters. In scenarios with continuous sequences of actions evaluated on Dataset UoG2020, RadMamba surpasses other models with significantly higher parameter counts by at least 3%, achieving this with only 6.7 k parameters. Our code is available at: https://github.com/lab-emi/AIRHAR.
Index Terms:
Human activity recognition, Mamba state-space model, Frequency Modulated Continuous Wave (FMCW) radar, micro-Doppler signatures, deep learning, parameter-efficient networks, continuous monitoring, radar signal processing
I Introduction
Human Activity Recognition (HAR) technologies have become increasingly critical in modern society, finding essential applications in healthcare, elderly care, smart homes, and security. Traditional HAR methods predominantly rely on wearable sensors or camera-based systems. While wearable devices offer precise monitoring, they often compromise user comfort and face challenges related to battery life and user compliance. Camera-based solutions, although versatile, introduce significant privacy concerns and are susceptible to performance degradation under challenging environmental conditions. Consequently, Radar-based Human Activity Recognition (RadHAR) systems have emerged as a compelling alternative due to their inherent advantages, including user privacy preservation, robustness to light conditions, non-intrusive and through-wall sensing capabilities [1, 2].
The field of radar-based HAR has evolved considerably from initial works focusing on simple scenarios with clearly defined activities [3, 4] toward more complex and realistic settings, including continuous activity monitoring [5, 6], multi-modal and distributed sensing [7, 8], and advanced classification techniques based on machine learning and deep learning methods. Early solutions predominantly relied on conventional machine learning algorithms, such as Support Vector Machines and k-Nearest Neighbors (kNN), utilizing handcrafted features like micro-Doppler centroid, bandwidth, and statistical measures [4, 5, 6]. Although these traditional techniques demonstrated acceptable performance in constrained scenarios, they encountered challenges related to scalability, complexity in feature engineering, and generalization to realistic continuous environments.
The advent of deep learning techniques, particularly convolutional neural networks (CNNs) [9, 10], recurrent neural networks (RNNs) [11, 7], and hybrid CNN-RNN models [12, 13, 14], revolutionized radar-based HAR by enabling automatic extraction and representation of complex spatial-temporal features directly from radar data. However, despite their accuracy, CNN and hybrid models typically suffer from high computational complexity and latency, limiting their suitability for real-time deployment, especially in resource-constrained edge scenarios.
Recently, advanced transformer-based architectures such as ViT and SSM have demonstrated significant improvements in feature modeling capabilities and explored lightweight designs across various domains [15, 16, 17, 18, 19, 20]. Nevertheless, their computational complexity remains high, making it challenging to balance the number of parameters and the number of floating-point operations per inference (#FLOP/Inf.). For instance, the FML-ViT model with only 169 M #FLOP/Inf. requires 2.7 M parameters [16], while model LW-ViT with 769 k parameters demands significantly higher #FLOP [21]. Compared to ultra-lightweight object detection models optimized for edge devices, such as YOLO-Fastest with 350 k parameters, and 252 M #FLOP/Inf, the direct application of these advanced architectures to radar-based HAR tasks remains costly due to their high computational overhead. This limitation is particularly critical for applications with limited resources, such as gesture recognition in assistive technologies on mobile or embedded devices [22]. Additionally, applications like smart home monitoring, which often involve multiple sensors deployed across a home, would greatly benefit from lightweight and energy-efficient algorithms [8].
To overcome these limitations and leverage the modeling strengths of state-space architectures while ensuring computational efficiency in radar-based HAR, we introduce a novel SSM Mamba-based architecture named RadMamba. In summary, the main contributions of this work are as follows:
  • We propose RadMamba, a parameter-efficient radar-based HAR architecture incorporating Channel Fusion and Downsampling, Doppler-aligned segmentation, and convolutional token projections, specifically tailored to address the unique characteristics of micro-Doppler radar signatures.
  • RadMamba achieves state-of-the-art classification accuracy on multiple radar-based HAR datasets: it matches the best accuracy of existing heavyweight models on non-continuous datasets (99.8% accuracy on Dataset DIAT and 91.2% accuracy on Dataset CI4R), and surpasses existing models by at least 3% on scenarios with continuous sequences of activities (89.3% accuracy on Dataset UoG2020), all while significantly reducing the number of parameters. Moreover, its computational complexity is the lowest compared to previous models validated on Dataset CI4R and Dataset UoG2020.
  • We perform a comprehensive ablation study demonstrating the critical role and synergy of each proposed architectural component, showing that the combination of micro-Doppler-aligned segmentation, adaptive downsampling, and convolutional projections is essential for optimal performance.
  • We provide an open-source PyTorch framework, RadHAR, facilitating rapid prototyping and evaluation of micro-Doppler-based HAR methods, promoting reproducibility and future research.
II Related Works
II-A Conventional machine learning method
These methods involved extracting domain-specific features from radar signals, followed by classification using algorithms like kNN and SVM. For instance, Erol et al. [5] used kNN to analyze power burst curves in spectrograms for fall detection, distinguishing fall from non-fall events. Li et al. [6] applied SVMs to micro-Doppler features such as Doppler centroid, bandwidth, and Singular Value Decomposition (SVD) components, while Ding et al. [23] extracted 28 features of four types: dynamic Doppler frequency, range change, energy change and dispersion of range and Doppler from dynamic range-Doppler trajectories for subspace kNN classification. Although effective in controlled environments, these approaches struggled to generalize across diverse conditions and activities due to their dependence on hand-crafted features, often leading to overfitting and poor scalability for continuous tasks [1, 2].
II-B CNN-based method
The shift to deep learning marked a turning point for RadHAR, with Convolutional Neural Networks leading the charge. Unlike traditional methods, CNNs automatically learn spatial hierarchies from radar representations like spectrograms and range-Doppler maps. Kang et al. [24] employed CNNs in a two-step process for segmenting and classifying activity sequences, achieving over 97% segmentation accuracy and 95% classification accuracy with 60 GHz millimeter-wave radar data. Yu et al. [10] introduced a ‘Dual-View CNN’ operating on orthogonal projections of voxelized point clouds, reaching 97.61% accuracy. However, CNNs applied solely on micro-Doppler maps simply treated as images, lacks inherent temporal modeling, a critical limitation for continuous activity recognition.
II-C RNN-based method
To address temporal dynamics in human activities, researchers incorporated Recurrent Neural Networks, particularly Long Short-Term Memory (LSTM[25] and Gated Recurrent Unit (GRU[26] networks, specifically designed for sequential data processing. Shrestha et al. [27] demonstrated the superiority of bidirectional LSTM networks over traditional SVMs and standard LSTMs in continuous activity recognition. Werthen-Brabants et al. [11] proposed an innovative split Bidirectional (Bi)-RNN architecture for real-time fall alerts, using separate forward & backward RNN components for immediate, refined predictions.
II-D Hybrid CNN-RNN-based method
Recognizing the complementary strengths of CNNs and RNNs, hybrid architectures emerged as a dominant paradigm. These models typically combine CNN-based spatial feature extraction with RNN-based temporal modeling. Kurtoglu et al. [12] implemented a sophisticated multi-modal approach, utilizing 3D CNNs with Bi-LSTM layers for range-Doppler maps and 2D/1D CNNs with Bi-LSTM for spectrograms and envelopes, achieving 93.3% accuracy for continuous activities mixed with American Sign Language [28]. Zhu et al. [13] integrated CNN-extracted features with LSTM-based temporal modeling, achieving 98.65% accuracy on a non-continuous Doppler radar human activity recognition task with 7 classes.
For distributed radar systems, hybrid models achieve an accuracy of 87.1% in continuous activity recognition tasks with only 71 k parameters [14]. Compared to CNN-based models in non-continuous tasks that utilize 135 M parameters [29] and RNN-based models in continuous tasks that utilize 189 M parameters [27] to achieve their best classification accuracy, hybrid models are more lightweight. However, the Floating-point Operation (FLOP) required by the hybrid model [14] is evaluated to be at least 1 G. The arithmetic intensity of CNNs, especially for multi-channel signals, drives their computational complexity.
II-E Transformer and SSM-based method
Recently, researchers have explored transformer-based architectures and SSM in RadHAR. ViT-based methods [16] achieved 92% accuracy with 2.7 M parameters, while [21] reached 92.1% with 769 k parameters on non-continuous FMCW radar data [30], though both demand significant ##\##FLOP per inference. (169 M and 2.41 G, respectively). In continuous HAR, however, ViT appears to under-perform conventional CNN-based models, such as the first to the third blocks of AlexNet and ResNet [17]. SSM-based approaches [19] hit 99.82% accuracy on a non-continuous Continuous Wave (CW) dataset [30] with 8.7 M parameters and 1.22G ##\##FLOP per inference.
III Analysis of Radar Data fed to ViM and Design Motivation
Figure 1: The schematic of conventional Vision Mamba and its three potential issues while classifying the micro-Doppler-based human activity recognition.
To effectively combine the modeling strength of SSM and decrease computational complexity, it is crucial to consider the unique properties of micro-Doppler signatures. To illustrate this, we trace the data flow from input to SSM in Vision Mamba (ViM), highlighting issues specific to radar micro-Doppler-based HAR tasks. Other details will be explained in section IV. As shown in Fig. 1, classical ViM comprises three core components: patch and position embedding, Mamba block with bidirectional branches, and SSMs.
III-A Issue 1: Micro-Doppler signatures exhibit significant sparsity.
The micro-Doppler signature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ serves as input to the classifier, with the Doppler frequency dimension represented as the image height $H$ and time as the image width $W$. The number of channels $C$ may correspond to channels from multiple radar sensors or one of the RGB channels. Analysis of Dataset [31] CI4R reveals that $\mathbf{X}$ exhibits an average spectrogram sparsity of approximately 87%, as illustrated in Fig.1 Issue 1. In the subsequent Segment step, $\mathbf{X}$ is divided into a matrix of smaller patches $\mathbf{X}_{seg}\in\mathbb{R}^{C\times I\times J\times H_{seg}\times W_{seg}}$, where $H_{seg}$ and $W_{seg}$ denote the height and width of each patch, respectively. Consequently, the number of patches per column is $I=\frac{H}{H_{seg}}$, and per row is $J=\frac{W}{W_{seg}}$. As depicted in Fig.1 Issue 1, many patches in $\mathbf{X}_{seg}$ contain only background information and resemble one another, providing little discriminative detail regarding Doppler frequency and time within each patch.
III-B Issue 2: Destruction of Spatial and Temporal Information.
Changes in Doppler frequency over time are critical for distinguishing various human activities. However, rectangular patch segmentation disrupts the relative positioning of Doppler frequency and time through two processes:
III-B1 Segment Undermines Temporal Interpretability in SSM, as Shown in Fig. 1, Issue 2 Left
As illustrated in Fig.1 left, a patch with column index $j$ can be interpreted as time point $t_{j}$, where $j=0,1,\cdots J$. During segmentation, the patch matrix of dimensions $I\times J$ is reshaped into a one-dimensional sequence $N=I\cdot J$. This reshaping results in a time point sequence of dimension $N$ as $[t_{0},t_{1},\cdots,t_{J},t_{0},t_{1},\cdots,t_{J},\cdots]$, represented by the yellow folded line in Fig.1 left. However, in a real-time sequence, $t_{J}$ and $t_{0}$ are not continuous. When the dimension $N$ remains unchanged from Segment to SSMbranch, and the SSM branches treat $N$ as the direction of temporal evolution, the SSM encounters a discontinuous sequence. This discontinuity hinders its ability to discern significant features in the micro-Doppler signature.
III-B2 Flatten Compromises Spatial Information Integrity Across Time and Doppler Dimensions in Patches, as Shown in Fig. 1, Issue 2 Right
Following segmentation, each patch is flattened into a vector:
$\mathbf{X^{\prime}_{seg}}=\textit{Flatten}(\mathbf{X_{seg}})$ (1)
where $\mathbf{X^{\prime}{seg}}\in\mathbb{R}^{N\times CH{seg}W_{seg}}$. A pixel with row index $h_{seg}$ can be interpreted as Doppler frequency point $f_{h_{seg}}$, where $h_{seg}=0,1,\cdots H_{seg}$. The Flatten operation results in a Doppler point sequence in $H_{seg}W_{seg}$ as $[f_{0},f_{1},\cdots,f_{h_{seg}},f_{0},f_{1},\cdots,f_{h_{seg}},\cdots]$, as shown in Fig. 1 left. However, $f_{h_{seg}}$ and $f_{0}$ are not continuous in the Doppler frequency sequence, disrupting the Doppler frequency features. This disruption also affects the time dimension within each patch.
In summary, the second issue encountered by conventional ViM is that Doppler and temporal information within $H_{seg}$ and $W_{seg}$ is compromised by flattening.
III-C Issue 3: Challenges for Projection Layers in Capturing Doppler Patterns
Within the Mamba block, as shown in Fig. 1 top, the input $\mathbf{X_{p}}$ first undergoes layer normalization to yield $\mathbf{X}_{proj}$, which is then split into two branches via linear projections P1 and P2:
$\mathbf{Z}$ $=\mathbf{W}_{\textit{P1}}\mathbf{X}_{proj}+\mathbf{b}_{\textit{P1}},$ (2)
$\mathbf{X}_{fw}$ $=\mathbf{W}_{\textit{P2}}\mathbf{X}_{proj}+\mathbf{b}_{\textit{P2}},$ (3)
where the $\mathbf{W}$ and $\mathbf{b}$ terms are the weight matrices and bias vectors, respectively.
The projection operations are expected to extract Doppler frequency patterns over time and enhance the average cross-correlation between patches. This enhancement occurs because a high-performing classifier should recognize all Doppler-aligned vectors of a specific activity as a cohesive and highly relevant unit. However, the averaged cross-correlation between patches shows limited improvement from conventional projections P1 and P2. For instance, in Dataset CI4R [31], the average patch cross-correlation is 0.008 before projections 1 and 2. After projections 1 and 2, they rise to 77.9 and 0.4, respectively. In contrast to the improvements achieved through our proposed Conv1D projection (detailed later in Section IV-C), the linear layers in the conventional ViM architecture demonstrate limited capability to enhance cross-correlation between patches and consistently identify all time bins as components of the same action.
The tensors $\mathbf{X}{fw}$ and $\mathbf{X}{bw}$ resulting from the projections P1 and P2 are then passed through SSM branches. The outputs from both directions are combined using a gating mechanism modulated by the SiLU-activation to produce the gated output $\mathbf{Y}{sum}$. This gated output passes through projection layer P3, which is linear in the conventional ViM, and is added to the input via a residual connection:
$\mathbf{Y}_{out}=\mathbf{W}_{\textit{P3}}\mathbf{Y}_{sum}+\mathbf{b}_{\textit{P3}}+\mathbf{X}_{proj},$ (4)
Before and after projection 3, again using Dataset CI4R as an example, the average patch cross-correlation is 160.1 and 75.5, respectively. Essentially, the averaged cross-correlation between patches is even decreased by the conventional projection layer 3, posing a challenge to achieve good classification performance.
In the analysis of Issue 3, to compute the averaged patch cross-correlation, we first trained a model to its optimal performance, then duplicated and froze its parameters. By re-feeding the data into the frozen model, we calculated the averaged patch cross-correlation during inference. This averaged patch cross-correlation also accounts for different batches (classes). For example, with $\mathbf{X}{proj}\in\mathbb{R}^{B\times N\times D}$, each patch is labeled with its index $n$ in $\mathbf{X}{seg}$. The averaged patch cross-correlation is defined as:
$Corr_{avg}=$ $\frac{1}{B}(\frac{1}{N}\sum_{n=1}^{N}(\frac{1}{N}\sum_{n^{\prime}=1}^{N}($
$corr(\mathbf{X_{proj}(b,n)},\mathbf{X_{proj}(b,n^{\prime})}))))$ (5)
where the correlation between two vectors of equal length is computed as:
$corr(\mathbf{A},\mathbf{B})=\sum_{m=0}^{M}\sum_{m^{\prime}=0}^{M}\mathbf{A}[m]\cdot\mathbf{B}[m+m^{\prime}]$ (6)
IV Proposed RadMamba
To overcome the aforementioned challenges and issues, we propose a micro-Doppler-oriented RadMamba architecture, illustrated in Fig. 2.
Figure 2: The schematic of the proposed RadMamba architecture and its novel solutions for radar micro-Doppler-based classification.
These modifications to the conventional approach include channel fusion and down-sampling, Doppler-aligned segmentation, and modified Mamba block rMamba.
IV-A Solution 1: Channel Fusion and Downsample
To tackle the issue of many similar and empty patches caused by high sparsity inherent in micro-Doppler signatures input $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, we introduce a preprocessing pipeline prior to the segmentation stage, as depicted in Fig. 2 Chan-DS: Channel Fusion and Downsample.
$\mathbf{X_{c-d}=\textit{Chan-DS}(\mathbf{X})}$ (7)
where $\mathbf{X}\in\mathbb{R}^{C_{c-d}\times H_{c-d}\times W_{c-d}},H_{c-d}<H,W_{c-d}<W$ is the input micro-Doppler signature after the channel fusion and downsample for each batch. The architecture of the Chan-DS block is as follows:
[…]
where $L=1,2$. The Two-Dimensional (2D) convolutional layer and batch normalization layer mix channel information and extract preliminary features from the input tensor $\mathbf{X}$. A 2D max pooling layer and a One-Dimensional (1D) max pooling layer reduces spatial dimensions. Optionally, an additional 1D average pooling layer with a kernel height as one and kernel width larger than 1, for example (1, 4), further downsamples the time dimension if the data sparsity along the time axis, influenced by the Pulse Repetition Interval (PRI) of the radar and pre-processing to generate the input, is excessive.
In combination with the Doppler-Aligned Segment explained in the next subsection IV-B, this method avoids the issue of having many similar and empty patches of data. Moreover, the down-sampling decreases the available size for patch segmentation, which enhances the energy-efficient design.
IV-B Solution 2: Doppler-Aligned Segment and Patch Embedding
To maintain the temporal and Doppler location information embedded in micro-Doppler signatures, we propose a new patch segmentation strategy, as shown in Fig. 2 Solution 2 Doppler-Aligned Segment, DA-Seg:
$\mathbf{X_{seg}}=\textit{DA-Seg}(\mathbf{X}_{c-d})$ (8)
Specifically, we define patches with a height of $H_{seg}=H_{c-d}$ (the full Doppler dimension) and a width of $W_{seg}=1$ (a single time bin of the spectrogram).
This approach ensures that the temporal evolution of state sequence $N$ remains as $[t_{0},t_{1},\cdots,t_{J}]$ and continuous as time sequence $W$. When the data is sent to the SSM, as depicted in Fig.2 Solution 2 left, this method avoids the loss of critical temporal information that conventional square patches of the ViM might face.
Moreover, this segmentation approach results in a pixel sequence dimension $CH_{seg}W_{seg}=H_{c-d}$ of $\mathbf{X}^{\prime}_{seg}$ that preserves the frequency pattern as $[f_{0},f_{1},\cdots,f_{h\_seg}]$, during Flatten operations within each Doppler bin, as shown in Fig.2 Solution 2 left, as:
$\mathbf{X^{\prime}_{seg}}=\textit{Flatten}(\mathbf{X_{seg}})=\mathbf{X_{seg}}$ (9)
IV-C Solution 3: Convolution Projections in rMamba Blocks
Finally, to increase the capability of projection layers to capture the Doppler patterns, we replace the conventional linear projection layers within the conventional Mamba block with 1D convolutional layers, as highlighted in Fig. 2 rMamba Block, in the blue boxes on the top-right part of the figure.
Within the rMamba block, the linear projections P1 and P2 are replaced by Conv1D layer as:
[…]
$\mathbf{X}_{bw}$ $=\text{flip state dim}(\mathbf{X}_{fw}),$ (12)
where $\mathbf{Z},\mathbf{X}_{fw},\mathbf{X}_{bw}\in\mathbb{R}^{N\times dim}$. Both Conv1D operations utilize a kernel size (in channels, out channels, kernel size) = $(dim,dim,3)$.
When $\mathbf{X}_{fw}$ and $\mathbf{X}_{bw}$ pass through the SSM branches with architecture of:
[…]
The outputs from both directions are combined using a gating mechanism modulated by the SiLU-activated $\mathbf{Z}$ and summed:
$\mathbf{Y^{\prime}}_{fw/bw}$ $=\mathbf{Y}_{fw/bw}\odot\text{SiLU}(\mathbf{Z})$ (14)
$\mathbf{Y}_{sum}$ $=\mathbf{Y^{\prime}}_{fw}+\mathbf{Y^{\prime}}_{bw}$ (15)
where $\odot$ denotes element-wise multiplication. Additionally, in our proposed rMamba, the projection layer P3 is also replaced by Conv1D, so that:
$\mathbf{Y}_{out}=\text{Conv1D}_{P3}(\mathbf{Y}_{sum})+\mathbf{X}_{proj},$ (16)
where the final Conv1D has weights of dimensions $(dim,dim,1)$ and transforms the combined sequence back to $(N,dim)$.
These convolutional layers enable the model to capture subtle variations in Doppler-span changes over time, which are crucial to describe the kinematic evolution of the different activities to be classified. Also, let us take Dataset CI4R as a comparison reference and follow the method to calculate cross-correlation as explained in Section III Issue 3. While passing through RadMamba, the average time bin cross-correlation is 0.002 before projection 1 and 2.
Then, compared to conventional linear projection, the convolution projection 1 increased the averaged cross correlation gap between projection input and output from 77.9 to 1065.7. The convolutional projection 2 increased this gap from 0.4 to 16.1. The convolutional projection 3 increased the average time cross-correlation from 245.3 to 356.4, while the linear projection 3 decreased this number. This enhancement significantly improves the model’s sensitivity to Doppler patterns between patches and makes it easier to regard all patches as one activity, thereby boosting classification performance.
IV-D Selective State-Space Modeling for Micro-Doppler Sequences
Figure 3: The architecture of State Space Model in RadMamba.
The SSM is a significant layer in rMamba enables efficient sequence modeling for the dynamic nature of micro-Doppler signatures. As shown in Fig. 3 (a), the discrete SSM is defined as:
$\mathbf{h}[n]$ $={\mathbf{\bar{A}}}[n]\mathbf{h}[n-1]+{\mathbf{\bar{B}}}[n]\mathbf{x}[n],$ (17)
$\mathbf{y}[n]$ $=\mathbf{C}\mathbf{h}[n]+\mathbf{D}\mathbf{x}[n],$ (18)
where the state transition ${\mathbf{\bar{A}}}$ and the input matrices dimension ${\mathbf{\bar{B}}}$ are discretized from corresponding matrices ${\mathbf{A}}$ and ${\mathbf{B}}$ in continuous SSM by a learnable step size $\Delta$ as shown in Fig. 3 (b):
$\bar{\mathbf{A}}$ $=\exp(\boldsymbol{\Delta}\cdot\mathbf{A}),$ (19)
[…]
The input of the SSM branches in the proposed rMamba block are tensors $\mathbf{X}_{fw/bw}\in\mathbb{R}^{N\times dim}$; after Conv1D and LayerNorm in SSM branch, this is termed as $\mathbf{X^{\prime}}_{fw/bw}\in\mathbb{R}^{N\times dim}$.
Therefore, the dimension of the discretized state transition ${\mathbf{\bar{A}}}$ is ${N\times dim\times{dim_{s}}}$ and the input matrices dimension ${\mathbf{\bar{B}}}$ is ${N\times N\times{dim_{s}}}$, respectively. While $N$ equals the time dimension in micro-Doppler signatures, the input matrices $\mathbf{\bar{B}}$ captures how the Doppler-aligned vector influences the state $\mathbf{h}[n]$, and the discretized state transition weight ${\mathbf{\bar{A}}}$ capture how the last state $\mathbf{h}[n-1]$. Moreover, the matrices $\mathbf{C}\in\mathbb{R}^{N\times dim_{s}}$ translate the current state $\mathbf{h}[n]$ to the output. The skip matrices $\mathbf{D}\in\mathbb{R}^{dim}$ add weighted input directly to output of SSM, which helps the recurrent part focus on the temporal evolution.
The selective mechanism in the SSM introduces input-dependent parameterization that enhances adaptability as shown in Fig. 3 (c). The parameters $\mathbf{B}$, $\mathbf{C}$, and $\boldsymbol{\Delta}$ are dynamically generated from the input sequence via linear projections:
$\mathbf{B}$ $=\mathbf{W}_{\mathbf{B}}\mathbf{X},$ (21)
$\mathbf{C}$ $=\mathbf{W}_{\mathbf{C}}\mathbf{X},$ (22)
$\boldsymbol{\Delta}$ $=\text{Softplus}(\mathbf{W}_{\boldsymbol{\Delta}}\mathbf{X}),$ (23)
[…]
IV-E Overall Architecture
To summarize the method, the proposed RadMamba pipeline in pseudo-code is presented in Algorithm I. The patch embedding in PEmb follows the method in  [15] and the position embedding method is a sinusoidal position encoding. [32].
Algorithm 1 Proposed RadMamba Pipeline
Input: Micro-Doppler Signature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$
Parameters: $dim$, $dim_{s}$, $dt\_rank$
Output: Classification Result $\mathbf{F}\in\mathbb{R}^{Q}$, where $Q$ is the number of classes
1:Channel Fusion and Downsample:
2:   $\mathbf{X_{c-d}}=\textit{Chan-DS}(\mathbf{X})$, $\mathbf{X}_{c-d}\in\mathbb{R}^{C_{c-d}\times H_{c-d}\times W_{c-d}}$
3:Doppler-Aligned Segmentation:
4:   $\mathbf{X_{seg}}=\textit{DA-Seg}(\mathbf{X}_{c-d})$, $\mathbf{X}_{seg}\in\mathbb{R}^{C_{c-d}\times N\times H_{c-d}}$
5:Patch and Position Embedding:
6:   $\mathbf{X^{\prime}}_{p}=\textit{PEmb}(\mathbf{X}_{seg})$, $\mathbf{X^{\prime}}_{p}\in\mathbb{R}^{N\times D}$
7:rMamba Block:
8:   $\mathbf{Y}_{out}=\textit{rMamba}(\mathbf{X}_{p})$, $\mathbf{Y}_{out}\in\mathbb{R}^{N\times dim}$
9:Classify:
10:   $\mathbf{F}=\mathbf{W}_{out}\mathbf{Y}_{out}$, $\mathbf{F}\in\mathbb{R}^{Q}$
V Open-source Datasets and Preprocessing for Radar-based HAR
To evaluate the efficacy of our proposed RadMamba framework, we utilize three distinct radar datasets. This section describes these datasets, namely the CW non-continuous Dataset DIAT, FMCW non-continuous Dataset CI4R, and FMCW continuous Dataset UoG2020, along with their respective pre-processing steps. More details on these datasets are summarized in Tab. I. Here the word ‘continuous’ refers to samples of data where sequences of human activities are performed one after the other in a continuous manner.
TABLE I: Information of Different Radar-based Activity Recognition Datasets Used in This Work.
Datasets DIAT CI4R UoG2020
Continuous No No Yes
Classes 6 11 6
Sensor X-band CW IWR1443 FMCW FMCW
Freq.(GHz) 10 77 5.8
Resolution 224$\times$224 224$\times$224 240$\times$1471
BW(MHz) - 750 400
Length per sample (s) 3 20 35
Train Samples 2646 528 39
Test Samples 759 132 6
Sliding Window 224
V-A Dataset DIAT: X-Band CW Radar with Discrete Activity Instances [30]
This dataset was collected using an X-band CW radar operating at a frequency of 10 GHz. It encompasses data from 30 human subjects performing ‘suspicious’ activities at ranges spanning 10 meters to $\pm$0.5 kilometers over 3-second durations, resulting in an imbalanced set of 3,780 spectrogram samples of shape $(3,244,244)$across six activity classes. Preprocessing adheres to the methodology outlined in [30]. To facilitate a fair comparison with previous works, the dataset is partitioned into 70% training, 10% validation, and 20% test sets within each class, maintaining class-specific balance for reliable performance evaluation.
V-B Dataset CI4R: Short-Range FMCW Radar for Multi-Class Activity Recognition [31]
The experiments of the second dataset employ a Texas Instruments IWR1443 FMCW radar, operating at 77 GHz with a bandwidth of 750 MHz. It includes data from six participants, varying in age, height, and weight, performing 11 distinct activities and ambulatory gaits. Each participant executed 10 repetitions per activity, yielding 60 samples per class per sensor, as detailed in [31]. Preprocessing involves applying a 256-point short-time Fourier transform (STFT) with a context window length to convert the raw radar signals into a 2D micro-Doppler representation of shape $(1,244,244)$. The dataset is split with an 8:2 training-to-test ratio, ensuring sufficient data for model optimization and evaluation.
V-C Dataset UoG2020: Continuous FMCW Radar for Realistic HAR Sequences [27]
Representing a leap in complexity, this dataset was acquired using an FMCW radar operating at 5.8 GHz with a 400 MHz bandwidth. It comprises data from 15 participants (14 males, 1 female, aged 21–35) performing six basic activities within continuous 35-second sequences. These activities are organized into three distinct sequential orders, with participants granted autonomy in selecting transition points between consecutive activities and determining trajectories for translational tasks, as described in [27]. The resulting micro-Doppler signatures are preprocessed into $(1,240,244)$ 2D frames, with two participants’ data designated as the test set, while the remaining 13 contribute to training, providing a challenging benchmark for continuous sequence modeling.
VI Experimental Setup
VI-A Training Hyperparameter
Training is conducted using the ADAMW optimizer [33, 34], paired with the ReduceLROnPlateau scheduler to adjust the learning rate based on test performance adaptively. For Dataset DIAT, the initial learning rate is set to $1\times 10^{-4}$, with a batch size of 16, and training proceeds for 50 epochs. For Dataset CI4R, RadMamba employs an initial learning rate of $5\times 10^{-3}$, while the baseline models use $1\times 10^{-4}$, with a batch size of 16 over 100 epochs, reflecting the dataset’s increased complexity and sample size. For Dataset UoG2020, the initial learning rate is $5\times 10^{-5}$, with a larger batch size of 256 and 50 epochs, processing continuous sequences with a frame length of 224 and a stride of 1 to capture temporal dependencies effectively. All experiments are executed on a single NVIDIA RTX 4090 GPU.
To account for training variability, all models, including RadMamba, are trained and evaluated ten times with distinct seeds, reporting in the following analysis the mean and standard deviation of performance metrics in the model size swept experiment.
Figure 4: Accuracy vs. number of parameters (1) and number of ##\##FLOP per inference (2) on: (a) Dataset DIAT (b) Dataset CI4R (c) Dataset UoG2020
VI-B Network Structure
TABLE II: Network Architecture of Models Used in This Work
Model Swept Hidden Size ($dim$ for RadMamba)
ResNet [5, 7, 12, 20, 30]
Bi-LSTM [1, 2, 4, 10, 20, 28, 34]
CNN-LSTM [2, 4, 6, 16, 28, 34, 46]
CNN-Bi-GRU [1, 2, 4, 6, 16, 28]
RadMamba Dataset DIAT [8, 16, 32, 64, 80]
RadMamba Dataset CI4R [8, 16, 32, 64, 80, 96, 128, 160]
RadMamba Dataset UoG2020 [8, 16, 20, 24, 32]
TABLE III: Specified Network Architecture Variations of RadMamba
Datasets $dim_{s}$ $dt\_rank$ Chan. Fusion DS Redu. Factor
Dataset DIAT 1 2 [2, (16, 3, 3)] (2, 2)
Dataset CI4R 4 0 [1, (1, 3, 3)] (2, 8)
Dataset UoG2020 16 4 [1, (1, 3, 3)] (2, 32)
The classification performance of RadMamba is evaluated against a suite of established models, selected for their proven efficacy across CNN, RNN, and hybrid architectures: ResNet [35], Bi-LSTM [27], CNN-LSTM [13], and CNN-Bi-GRU [14]. To explore the impact of model scaling, we conducted a model size swept experiment, varying the hidden dimensions as shown in Tab. III. These sizes are chosen to capture the full spectrum of performance trends as model capacity increases, providing insights into scalability and efficiency.
Notably, the CNN-Bi-GRU model, originally designed for distributed radar systems in  [14], exhibits a larger parameter count, though its channel fusion approach may not generalize across all datasets.
For RadMamba, specific configurations of the channel fusion and down-sampling blocks, as well as the SSM parameters, are tailored to each dataset to optimize micro-Doppler processing as shown in Tab. III, where the ‘channel fusion’ column specifies [L, ($C_{c-d},H_{k},W_{k}$)], where L is the number of channel fusion blocks and the $C_{c-d},H_{k},W_{k}$ are channel, kernel height, kernel width of Conv2D layer, respectively. The ‘DS Redu. Factor’ denotes the reduction factors ($H/H_{c-d}$, $W/W_{c-d}$).
For transformer-based and SSM-based models such as FML-ViT [16], LW-ViT [21], and ActivityMamba [19], we adopt their reported results as optimal outcomes under minimal model sizes, as dimension definitions may vary and are not always fully specified in the original works.
VII Results and Discussion
VII-A Model Scaling and Efficiency Analysis
The model size swept results, illustrated in Figure 4, compare accuracy against the number of parameters from 2k to 150k and ##\##FLOP per inference across Datasets DIAT, CI4R, and UoG2020 between previous models ResNet, Bi-LSTM, CNN-Bi-GRU, and CNN-LSTM and the proposed RadMamba.
For Dataset DIAT, as presented in Fig. 4 (a1) and (a2), RadMamba achieves a peak accuracy of 99.8% with only 21.7 k parameters. However, RadMamba overfits with a model size larger than 21.7 k because the dataset is very simple and the accuracy is close to 100%. In contrast, ResNet’s accuracy increases steadily with model size, reaching 99.9% with 208.1 k parameters, yet it requires nearly ten times the parameters of RadMamba for slightly lower performance.
TABLE IV: Mean and Standard Deviation of Classification Accuracy Performance across Seed 0 to 9 of Different NN-based RadHAR Models Evaluated With Three Different Radar Settings alongside Their Model Size and Floating-point Operations per Inference Sample (FLOP/Inf.)
Dataset DIAT [30]
(CW, Non-continuous)
Dataset CI4R [31]
(FMCW, Non-continuous)
Dataset UoG2020 [27]
(FMCW, Continuous)
Classifiers
#params
#FLOP/Inf. a
Accuracy
#params
#FLOP/Inf. a
Accuracy
#params
#FLOP/Inf. a
Accuracy
ResNetb [35] 69.4 253.3 99.8$\pm$0.16 150.7 310.2 32.3$\pm$3.35 24.7 72.8 83.6$\pm$1.73
Bi-LSTM [27] 23.3 5.6 93.9$\pm$0.68 134.4 6.0 79.4$\pm$1.23 15.5 2.1 85.3$\pm$2.01
CNN-LSTM [13] 115.3 8.7 98.4$\pm$0.33 96.5 8.7 86.1$\pm$1.32 10.3 1.2 86.1$\pm$2.22
CNN-Bi-GRU [14] 55.0 1090.1 98.5$\pm$0.41 142.8 1013.0 84.9$\pm$1.11 53.6 1082.0 85.4$\pm$2.19
FML-ViTc [16] - - - $\sim$2700 169.0 $\sim$92.0 - - -
LW-VITc [21] - - - 769 $\sim$2410 $\sim$92.0 - - -
ActivityMambac [19] $\sim$8700.0 $\sim$1220.0 99.8 - - - - -
RadMamba (Ours) 21.7 145.6 99.8$\pm$0.12 71.4 8.8 91.2$\pm$2.57 6.7 1.0 89.3$\pm$1.38
RadMamba (Ours) 10.1 14.5 98.2$\pm$0.46 47.9 6.0 88.8$\pm$4.58 17.7 1.8 89.6$\pm$0.78
  • a
    ##\##FLOP/Inf. is the number of Floating point operations per inference.
  • b
    This ResNet only retains the structure of ResNet18 from the first layer to the end of the second residual block. The relative multiples of channels between convolution layers remain unchanged, but the size of the minimum channel can be adjusted.
  • c
    The $\sim$’ indicates that these results are approximated from the references [16, 21, 19].
Figure 5: Comparison between ground truth and best prediction of RadMamba model with the configuration in Table I on (a) Dataset DIAT; (b) Dataset CI4R; (c) Dataset UoG2020.
On Dataset CI4R, as shown in Fig. 4 (b1) and (b2), RadMamba consistently outperforms previous models across a range of parameter configurations in efficiency, maintaining superior accuracy with fewer parameters and ##\##FLOP/Inf. Similarly, for Dataset UoG2020 shown in Fig. 4 (c1) and (c2), RadMamba still records a higher accuracy for continuous activities scenarios with fewer ##\##FLOP/Inf., outperforming larger and more computationally intensive models in efficiency and effectiveness.
These trends, as depicted in the figure, affirm RadMamba’s capability to achieve superior accuracy across diverse radar modalities while drastically reducing both parameter count and ##\##FLOP/Inf., positioning it as an optimal solution for lightweight, real-time human activity recognition.
VII-B Comparison with Previous Works
Figure 6: Averaged test loss across 10 random seeds vs. Epochs during training of models with configuration in Table I on (a) Dataset DIAT; (b) Dataset CI4R; (c) Dataset UoG2020.
The experimental results detailed in Table IV highlight the superior performance of our proposed RadMamba model across multiple radar-based human activity recognition datasets.
On Dataset DIAT, which utilizes CW radar in a non-continuous setting, RadMamba achieved an accuracy of 99.8% with a standard deviation of $\pm$0.12%, using only 21.7k parameters and requiring 145.6 million ##\##FLOP/Inf. per sample. This demonstrates its capability to match the state-of-the-art accuracy of previous models while being significantly more parameter-efficient compared to $\sim$8700k parameters and computationally lighter compared to $\sim$1220M ##\##FLOP per inference. The CNN-LSTM with 8.7 M ##\##FLOP per inference achieves better accuracy of 98.4% than 98.2% from RadMamba with 14.5 M ##\##FLOP per inference. However, take 45nm technology for example, the energy consumption for 8kB SRAM memory access of one 32-bit floating point parameter and one 32-bit floating point computing operation is 5.0 pJ and 4.6 pJ [36]. With 11$\times$ larger model and around half ##\##FLOP per inference, it’s hard for CNN-LSTM to be more energy-efficient than RadMamba.
TABLE V: Ablation Study with progressive techniques (denoted by Tech.) of New Blocks in RadMamba across Seed 0 to 5, Evaluated with Three Different Radar Settings alongside Their Number of Parameters
Dataset DIAT Dataset CI4R Dataset UoG2020
Row Tech. 1: Proj. Layer Tech. 2: Patch Size
Tech. 3:
Down-sample.
Accuracy
#Params
Tech 3:
Down-sample.
Accuracy
#Params
Tech 3:
Down-sample.
Accuracy
#Param
1 (Conventional) (1, 1) 93.5 $\pm$1.06 11.3 (1, 1) 61.9 $\pm$2.55 40.6 (1, 1) 71.2 $\pm$2.10 4.2
[…]
3 ($H_{seg}$, $W_{seg}$)a (2, 2) 99.6 $\pm$0.06 11.3 $\Leftrightarrow$ (2, 8) 63.0 $\pm$7.75 40.6 $\Leftrightarrow$ (2, 32) 66.9 $\pm$2.01 4.2
4 (1, 1) 87.7 $\pm$1.25 39.3 (1, 1) 67.9 $\pm$0.76 54.9 (1, 1) 72.6 $\pm$2.20 7.6
[…]
6 (1, $W_{c-d}$) (2, 2) 98.5 $\pm$0.43 21.4 $\Leftrightarrow$ (2, 8) 77.3 $\pm$2.80 38.9 $\Leftrightarrow$ (2, 32) 64.0 $\pm$1.26 3.6
7 (1, 1) 96.9 $\pm$0.52 39.3 (1, 1) 88.4 $\pm$0.76 54.9 (1, 1) 83.9 $\pm$1.85 7.8
[…]
9 1 linear layer ($H_{c-d}$, 1) (2, 2) 99.9 $\pm$0.00 21.4 $\Leftrightarrow$ (2, 8) 86.5 $\pm$4.35 45.8 $\Leftrightarrow$ (2, 32) 84.2 $\pm$2.00 5.7
10 (1, 1) 92.6 $\pm$0.98 11.6 (1, 1) 40.1 11.98 66.2 (1, 1) 70.6 $\pm$2.12 5.2
[…]
12 ($H_{seg}$, $W_{seg}$)a (2, 2) 99.5 $\pm$0.10 11.6 $\Leftrightarrow$ (2, 8) 57.8 $\pm$9.64 66.2 $\Leftrightarrow$ (2, 32) 68.5 $\pm$1.50 5.2
13 (1, 1) 88.8 $\pm$2.42 39.6 (1, 1) 48.3 $\pm$9.80 80.5 (1, 1) 70.8 $\pm$2.43 8.6
[…]
15 (1, $W_{c-d}$) (2, 2) 98.7 $\pm$0.32 21.7 $\Leftrightarrow$ (2, 8) 72.4 $\pm$3.12 64.5 $\Leftrightarrow$ (2, 32) 64.0 $\pm$2.16 4.7
16 (1, 1) 97.2 $\pm$0.26 39.6 (1, 1) 85.3 $\pm$1.27 80.5 (1, 1) 83.6 $\pm$1.04 8.9
[…]
18 3 linear layerb ($H_{c-d}$, 1) (2, 2) 99.8 $\pm$0.06 21.7 $\Leftrightarrow$ (2, 8) 85.3 $\pm$2.67 71.4 $\Leftrightarrow$ (2, 32) 85.2 $\pm$1.16 6.7
19 (1, 1) 93.8 $\pm$1.14 11.6 (1, 1) 65.5 $\pm$3.30 66.2 (1, 1) 75.3 $\pm$2.34 5.2
[…]
21 ($H_{seg}$, $W_{seg}$)a (2, 2) 99.6 $\pm$0.15 11.6 $\Leftrightarrow$ (2, 8) 70.7 $\pm$3.23 66.2 $\Leftrightarrow$ (2, 32) 74.3 $\pm$2.31 5.2
22 (1, 1) 90.1 $\pm$1.72 39.6 (1, 1) 70.7 $\pm$3.43 80.5 (1, 1) 76.4 $\pm$2.77 8.6
[…]
24 (1, $W_{c-d}$) (2, 2) 98.9 $\pm$0.13 21.7 $\Leftrightarrow$ (2, 8) 79.0 $\pm$2.42 64.5 $\Leftrightarrow$ (2, 32) 68.8 $\pm$2.08 4.7
25 (1, 1) 96.8 $\pm$0.86 39.6 (1, 1) 87.9 $\pm$0.36 80.5 (1, 1) 84.5 $\pm$0.77 8.9
[…]
27 (Ours) 1 conv1d layerc ($H_{c-d}$, 1) (2, 2) 99.8 $\pm$0.11 21.7 $\Leftrightarrow$ (2, 8) 89.6 $\pm$3.69 71.4 $\Leftrightarrow$ (2, 32) 88.6 $\pm$1.46 6.7
  • a
    $(H_{1},W_{1})$ equals to (7, 7) for Dataset DIAT and B, (5, 7) for Dataset UoG2020 due to the different image size of these datasets.
  • b
    This option replaces projections 1 and 2 from a single linear layer to three linear layers.
  • c
    The kernel size for projection 1 and 2 is $(dim,3,3)$, and for projection 3 is $(dim,1,1)$.
For Dataset CI4R, employing FMCW radar in a non-continuous scenario with 11 activities, RadMamba attained an accuracy of 91.2% with a standard deviation of $\pm$2.57%, leveraging 71.4 thousand parameters. This performance outperforms previous models shown in Tab. IV. Moreover, it remains competitive with the state-of-the-art of previous models while requiring substantially fewer computational resources (8.8 M ##\##FLOP/Inf. vs. 169 M to 2410 M ##\##FLOP/Inf.), showcasing its ability to excel in more complex activity recognition tasks with minimal computational overhead. Moreover, compared to the Bi-LSTM model with lowest arithmetic intensity in table, RadMamba outperforms Bi-LSTM accuracy of 9.4% with 86.5 k less parameters and same ##\##FLOP per inference of 6.0 M.
On Dataset UoG2020, which involves FMCW radar in a continuous setting with 6 activities, RadMamba recorded an accuracy of 89.3% with a standard deviation of $\pm$1.38%, utilizing 6.7 thousand parameters and 1.3M ##\##FLOP per inference. With an even smaller model size and more complicated task compared to Datasets DIAT and CI4R, RadMamba achieves the best accuracy in models shown in Tab. IV and highlights its effectiveness in continuous HAR scenarios with the smallest model size and ##\##FLOP per inference.
Figure 5 (a) and (b) present confusion matrices comparing the ground truth labels with the best predictions of RadMamba, illustrating its classification accuracy and error distribution for each dataset. Figure 5 (c) shows the comparison between the classification results of RadMamba and CNN-LSTM over time. Complementing this, Figure 6 displays the averaged test loss across ten random seeds as a function of training epochs, providing insight into the model’s convergence behavior and stability during training on the same datasets.
VII-C Ablation study of the Micro-doppler-oriented Structure
Table V presents an ablation study evaluating the impact of new architectural blocks in the RadMamba model across three datasets with varying radar settings. The study systematically assesses the effects of different projection layers, patch sizes, and down-sampling scales on model accuracy and parameter efficiency, with results averaged over seeds 0 to 5.
At the outset, the most important conclusion is that optimal performance emerges from the combined integration of all three proposed new blocks—projection layer, patch size, and down-sampling scale—rather than any single component in isolation. This is evident in the incremental improvements observed across the configurations. For instance, the baseline conventional ViM model (Row 1) achieves an accuracy of 93.5% on Dataset DIAT, 61.9% on Dataset CI4R, and 71.2% on Dataset UoG2020. In contrast, the proposed RadMamba (Row 27), which incorporates a 1D convolution layer, a patch size of ($H_{c-d}$, 1), and a down-sampling (2, 2), achieves a higher accuracy of 99.8% on Dataset DIAT. Similarly, on Dataset CI4R, Row 27 with a (2, 8) down-sampling scale yields significantly higher accuracy of 89.6%, and on Dataset UoG2020 with a (2, 32) scale, it reaches 88.6%. Intermediate configurations, such as Row 19 (only projection replacement), Row 7 (only patch size changing) and Row 2 (only down-sampling applied) show lower accuracies—underscoring that no single structure alone achieves the peak performance of the fully combined design in Row 27.
To provide a more detailed analysis, we discuss the three blocks—projection layer, patch size, and down-sampling scale—separately, evaluating their individual contributions to RadMamba’s performance across the datasets.
VII-C1 Projection Replacement
The adoption of a 1D convolutional layer as the projection mechanism consistently outperforms configurations with a single linear layer or three linear layers across nearly all settings. Comparing Rows 19-27 with Rows 10-18 and Rows 1-9, the conv1d projection replacement within the same patch size and down-sampling scale achieved average increment of mean accuracy by 0.48% over 1 linear layer, 0.40% over 3 linear layer on Dataset DIAT. On Dataset CI4R and UoG2020, the conv1d layer shows a substantial improvement of 2.99% and 3.64% over 1 linear layer, respectively.
VII-C2 Patch Size
The ($H_{c-d}$, 1) configuration consistently delivers superior performance within each projection layer type. For the 1D convolutional layer, Row 25 with ($H_{c-d}$, 1) and without down-sampling achieves 3% gain on Dataset DIAT, 22.4% gain on Dataset CI4R, 9.2% gain on Dataset UoG2020 vs. Row 19 with rectangular patch. This trend holds across other projection types, indicating that the ($H_{c-d}$, 1) patch size effectively retains the critical spatial features for radar data processing.
VII-C3 Down-sampling
The results suggest that down-sampling along the patch vector dimension generally outperforms down-sampling along the dimension orthogonal to patch vector and no down-sampling. For instance, Row 27’s (2, 32) time down-sampling achieves 88.6% $\pm$ 1.46, outperforming Row 26’s (8, 2) at 84.4% $\pm$ 2.05 and Row 25’s (1, 1) at 84.5% $\pm$ 0.77 on Dataset UoG2020. This pattern highlights the advantage of downsampling in enhancing the micro-Doppler features.
The ablation study underscores the efficacy of the proposed RadMamba design, with the highest performance achieved by integrating a 1D convolutional projection layer, a ($H_{c-d}$, 1) patch size, and time-dimension down-sampling (Row 27).
VIII Conclusion
RadMamba sets a new benchmark in RadHAR, delivering top-tier accuracy with minimal parameters and computational cost across diverse datasets. Its radar micro-Doppler-oriented design addresses domain-specific challenges, making it ideal for real-time, resource-constrained applications. Overall, RadMamba consistently outperforms previous neural network models like ResNet, Bi-LSTM, CNN-LSTM, and CNN-Bi-GRU, and competes with or exceeds advanced architectures like Vision Transformers and ActivityMamba, all while maintaining a remarkably low parameter count (ranging from 6.7 k to 71.4 k) and reduced computational complexity (1.3 M to 145.6 M ##\##FLOP/Inf.). These results emphasize RadMamba’s capability as a lightweight and high-performance solution for radar-based HAR across diverse radar modalities and activity recognition tasks. The micro-Doppler-oriented design, integrating channel fusion, Doppler-aligned segmentation, and convolutional projections, addresses the unique challenges of radar data, delivering high accuracy across CW and FMCW datasets, both continuous and non-continuous. Ablation studies confirm the synergy of its components, with no single element achieving comparable performance alone.
Acknowledgments
This work was partially supported by the European Research Executive Agency (REA) under the Marie Skłodowska-Curie Actions (MSCA) Postdoctoral Fellowship program, Grant No. 101107534 (AIRHAR). We thank Prof. Dr. L.C.N. de Vreede and Prof. Dr. Alexander Yarovoy for their support.

### v3

@@
-RadMamba: Efficient Human Activity Recognition through Radar-based Micro-Doppler-Oriented Mamba State-Space Model
+RadMamba: Efficient Human Activity Recognition through a Radar-based Micro-Doppler-Oriented Mamba State-Space Model
 Abstract
-Radar-based Human Activity Recognition (HAR) has emerged as a promising alternative to conventional monitoring approaches, such as wearable devices and camera-based systems, due to its unique privacy preservation and robustness advantages.
-However, existing solutions based on convolutional and recurrent neural networks, although effective, are computationally demanding during deployment.
-This limits their applicability in scenarios with constrained resources or those requiring multiple sensors.
-Advanced architectures, such as Vision Transformer (ViT) and State Space Model (SSM) architectures, offer improved modeling capabilities and have made efforts toward lightweight designs.
-However, their computational complexity remains relatively high.
-To leverage the strengths of transformer architectures while simultaneously enhancing accuracy and reducing computational complexity, this paper introduces RadMamba, a parameter-efficient, radar micro-Doppler-oriented Mamba SSM specifically tailored for radar-based HAR.
-Across three diverse datasets, RadMamba matches the top-performing previous model’s 99.8% classification accuracy on Dataset DIAT with only 1/400 of its parameters and equals the leading models’ 92.0% accuracy on Dataset CI4R with merely 1/10 of their parameters.
-In scenarios with continuous sequences of actions evaluated on Dataset UoG2020, RadMamba surpasses other models with significantly higher parameter counts by at least 3%, achieving this with only 6.7 k parameters.
-Our code is available at: https://github.com/lab-emi/AIRHAR.
-Index Terms:
-Human activity recognition, Mamba state-space model, Frequency Modulated Continuous Wave (FMCW) radar, micro-Doppler signatures, deep learning, parameter-efficient networks, continuous monitoring, radar signal processing
+Radar-based Human Activity Recognition (HAR) is an attractive alternative to wearables and cameras because it preserves privacy, and is contactless and robust to occlusions.
+However, dominant Convolutional Neural Network (CNN)- and Recurrent Neural Network (RNN)-based solutions are computationally intensive at deployment, and recent lightweight Vision Transformer (ViT) and State Space Model (SSM) variants still exhibit substantial complexity.
+In this paper, we present RadMamba, a parameter-efficient, micro-Doppler–oriented Mamba SSM tailored to radar HAR under on-sensor compute, latency, and energy constraints typical of distributed radar systems.
+RadMamba combines (i) channel fusion with downsampling, (ii) Doppler-aligned segmentation that preserves the physical continuity of Doppler over time, and (iii) convolutional token projections that better capture Doppler-span variations, thereby retaining temporal–Doppler structure while reducing the number of Floating-point Operations/Inference (#\#FLOP/Inf.).
+Evaluated across three datasets with different radars and types of activities, RadMamba matches the prior best $99.8\%$ accuracy of a recent SSM-based model on the Continuous Wave (CW) radar dataset, while requiring only $1/400$ of its parameters.
+On a dataset of non-continuous activities with Frequency Modulated Continuous Wave (FMCW) radar, RadMamba remains competitive with leading $92.0\%$ results using about $1/10$ of the parameters, and on a continuous FMCW radar dataset it surpasses methods with far more parameters by at least $3\%$, using only $6.7\,\mathrm{k}$ parameters.
+Code: https://github.com/lab-emi/AIRHAR.
 I Introduction
-Human Activity Recognition (HAR) technologies have become increasingly critical in modern society, finding essential applications in healthcare, elderly care, smart homes, and security.
-Traditional HAR methods predominantly rely on wearable sensors or camera-based systems.
-While wearable devices offer precise monitoring, they often compromise user comfort and face challenges related to battery life and user compliance.
-Camera-based solutions, although versatile, introduce significant privacy concerns and are susceptible to performance degradation under challenging environmental conditions.
-Consequently, Radar-based Human Activity Recognition (RadHAR) systems have emerged as a compelling alternative due to their inherent advantages, including user privacy preservation, robustness to light conditions, non-intrusive and through-wall sensing capabilities [1, 2].
-The field of radar-based HAR has evolved considerably from initial works focusing on simple scenarios with clearly defined activities [3, 4] toward more complex and realistic settings, including continuous activity monitoring [5, 6], multi-modal and distributed sensing [7, 8], and advanced classification techniques based on machine learning and deep learning methods.
-Early solutions predominantly relied on conventional machine learning algorithms, such as Support Vector Machines and k-Nearest Neighbors (kNN), utilizing handcrafted features like micro-Doppler centroid, bandwidth, and statistical measures [4, 5, 6].
-Although these traditional techniques demonstrated acceptable performance in constrained scenarios, they encountered challenges related to scalability, complexity in feature engineering, and generalization to realistic continuous environments.
-The advent of deep learning techniques, particularly convolutional neural networks (CNNs) [9, 10], recurrent neural networks (RNNs) [11, 7], and hybrid CNN-RNN models [12, 13, 14], revolutionized radar-based HAR by enabling automatic extraction and representation of complex spatial-temporal features directly from radar data.
-However, despite their accuracy, CNN and hybrid models typically suffer from high computational complexity and latency, limiting their suitability for real-time deployment, especially in resource-constrained edge scenarios.
-Recently, advanced transformer-based architectures such as ViT and SSM have demonstrated significant improvements in feature modeling capabilities and explored lightweight designs across various domains [15, 16, 17, 18, 19, 20].
-Nevertheless, their computational complexity remains high, making it challenging to balance the number of parameters and the number of floating-point operations per inference (#FLOP/Inf.).
-For instance, the FML-ViT model with only 169 M #FLOP/Inf.
-requires 2.7 M parameters [16], while model LW-ViT with 769 k parameters demands significantly higher #FLOP [21].
-Compared to ultra-lightweight object detection models optimized for edge devices, such as YOLO-Fastest with 350 k parameters, and 252 M #FLOP/Inf, the direct application of these advanced architectures to radar-based HAR tasks remains costly due to their high computational overhead.
-This limitation is particularly critical for applications with limited resources, such as gesture recognition in assistive technologies on mobile or embedded devices [22].
-Additionally, applications like smart home monitoring, which often involve multiple sensors deployed across a home, would greatly benefit from lightweight and energy-efficient algorithms [8].
-To overcome these limitations and leverage the modeling strengths of state-space architectures while ensuring computational efficiency in radar-based HAR, we introduce a novel SSM Mamba-based architecture named RadMamba.
-In summary, the main contributions of this work are as follows:
-  • We propose RadMamba, a parameter-efficient radar-based HAR architecture incorporating Channel Fusion and Downsampling, Doppler-aligned segmentation, and convolutional token projections, specifically tailored to address the unique characteristics of micro-Doppler radar signatures.
-  • RadMamba achieves state-of-the-art classification accuracy on multiple radar-based HAR datasets: it matches the best accuracy of existing heavyweight models on non-continuous datasets (99.8% accuracy on Dataset DIAT and 91.2% accuracy on Dataset CI4R), and surpasses existing models by at least 3% on scenarios with continuous sequences of activities (89.3% accuracy on Dataset UoG2020), all while significantly reducing the number of parameters.
-Moreover, its computational complexity is the lowest compared to previous models validated on Dataset CI4R and Dataset UoG2020.
-  • We perform a comprehensive ablation study demonstrating the critical role and synergy of each proposed architectural component, showing that the combination of micro-Doppler-aligned segmentation, adaptive downsampling, and convolutional projections is essential for optimal performance.
-  • We provide an open-source PyTorch framework, RadHAR, facilitating rapid prototyping and evaluation of micro-Doppler-based HAR methods, promoting reproducibility and future research.
+Human Activity Recognition (HAR) technologies are central to healthcare, elderly care, smart homes, and security.
+Traditional HAR solutions rely on wearable sensors or cameras.
+While wearables can offer precise monitoring, they suffer from issues related to user compliance, battery life, and comfort.
+Camera-based approaches raise privacy concerns and are at high risk of degradation under challenging environmental conditions.
+Consequently, Radar-based Human Activity Recognition (RadHAR) has emerged as a compelling alternative, leveraging active electromagnetic sensing to infer behaviors while preserving privacy, maintaining robustness to illumination and occlusions, and enabling through-wall sensing [1, 2].
+Early RadHAR studies progressed from simple, isolated activities [3, 4] to more realistic settings involving continuous activity monitoring [5, 6], multi-modal and distributed sensing [7, 8], and advanced classifiers.
+Conventional machine learning with hand-crafted micro-Doppler features (e.g., centroid, bandwidth, statistical descriptors, and Singular Value Decomposition (SVD) components) [4, 5, 6, 9] demonstrated proof of concept but struggled with scalability and generalization in continuous scenarios.
+Deep learning improved accuracy by learning spatio-temporal representations directly from radar signatures.
+CNN-based methods [10, 11, 12, 13], RNN-based methods [14, 7], and hybrid CNN-RNN models [15, 16, 17] substantially advanced performance.
+However, they often have high computational cost and latency, which limits on-sensor deployment in resource-constrained radar nodes.
+More recently, transformer and SSM-based architectures [18, 19, 20, 21, 22, 23] have shown strong modeling capacity and explored lightweight designs, but balancing parameter count with computational efficiency, measured by the number of floating-point operations per inference (#FLOP/Inf.), remains challenging for mobile and distributed applications such as gesture interfaces [24] and smart-home radar networks [8].
+Micro-Doppler signatures encode the evolution of Doppler frequency over time.
+Preserving this physical continuity is crucial for maintaining the high accuracy of sequential models; however, common vision-style rectangular patching and tokenization can disrupt Doppler–time ordering and waste computational resources on sparse, low-information regions.
+Moreover, edge radar deployments impose strict limits on memory, latency, and energy per inference.
+These constraints motivate architectures that (i) respect the physical features of radar signatures and sparsity patterns and (ii) reduce arithmetic intensity without sacrificing discriminative Doppler structure.
+In this work, we propose RadMamba, a radar-oriented, parameter-efficient Mamba SSM for micro-Doppler classification.
+RadMamba replaces vision-style patching and purely linear token projections with three radar-driven mechanisms: (i) channel fusion with downsampling to mix channels and shrink low-information regions; (ii) Doppler-aligned segmentation that preserves Doppler–time continuity for state-space modeling; and (iii) convolutional token projections to capture Doppler-span variations better and improve inter-patch coherence for the SSM backbone.
+The main contributions of this work are:
+  • A radar-centric state-space architecture that preserves the physical structure of micro-Doppler signatures via Doppler-aligned segmentation and convolutional token projections, reducing #\#FLOP/Inf.
+while maintaining temporal–Doppler fidelity.
+  • A comprehensive evaluation across CW and FMCW datasets covering non-continuous and continuous settings: RadMamba attains $99.8\%$ on a dataset measured by the Defence Institute of Advanced categorization problems and classification Technology, India (DIAT) [25], $91.2\%$ on a dataset from the laboratory of Computational Intelligence for Radar (CI4R) [26] (competitive with heavier models), and $89.3\%$ on a dataset collected at the University of Glasgow in 2020 (UoG20) [27] (at least $3\%$ higher than prior methods with far more parameters), with substantially fewer parameters and lower #\#FLOP/Inf.
+across all three datasets.
+  • An ablation study showing the synergistic effect of (i) channel fusion with downsampling, (ii) Doppler-aligned segmentation, and (iii) convolutional projections, each contributing to efficiency and accuracy beyond conventional ViT/SSM patching and linear projections.
+  • An open-source PyTorch framework, RadHAR, to facilitate reproducible research and rapid prototyping for micro-Doppler HAR.
 II Related Works
 II-A Conventional machine learning method
-These methods involved extracting domain-specific features from radar signals, followed by classification using algorithms like kNN and SVM.
+These methods involved extracting domain-specific features from radar signals, followed by classification using algorithms like k-Nearest Neighbors (kNN) and Support Vector Machine (SVM).
 For instance, Erol et al.
@@
 Li et al.
-[6] applied SVMs to micro-Doppler features such as Doppler centroid, bandwidth, and Singular Value Decomposition (SVD) components, while Ding et al.
-[23] extracted 28 features of four types: dynamic Doppler frequency, range change, energy change and dispersion of range and Doppler from dynamic range-Doppler trajectories for subspace kNN classification.
+[6] applied SVMs to micro-Doppler features such as Doppler centroid, bandwidth, and SVD components, while Ding et al.
+[9] extracted 28 features of four types: dynamic Doppler frequency, range change, energy change, and dispersion of range and Doppler from dynamic range-Doppler trajectories for subspace kNN classification.
 Although effective in controlled environments, these approaches struggled to generalize across diverse conditions and activities due to their dependence on hand-crafted features, often leading to overfitting and poor scalability for continuous tasks [1, 2].
 II-B CNN-based method
-The shift to deep learning marked a turning point for RadHAR, with Convolutional Neural Networks leading the charge.
+The shift to deep learning marked a turning point for RadHAR, with CNNs leading the charge.
 Unlike traditional methods, CNNs automatically learn spatial hierarchies from radar representations like spectrograms and range-Doppler maps.
@@
 [13] introduced a ‘Dual-View CNN’ operating on orthogonal projections of voxelized point clouds, reaching 97.61% accuracy.
-However, CNNs applied solely on micro-Doppler maps simply treated as images, lacks inherent temporal modeling, a critical limitation for continuous activity recognition.
+However, when CNNs are applied only to micro-Doppler maps treated as images, they lack inherent temporal modeling, which is a critical limitation for continuous activity recognition.
 II-C RNN-based method
-To address temporal dynamics in human activities, researchers incorporated Recurrent Neural Networks, particularly Long Short-Term Memory (LSTM[25] and Gated Recurrent Unit (GRU[26] networks, specifically designed for sequential data processing.
+To address temporal dynamics in human activities, researchers have incorporated RNNs, particularly Long Short-Term Memory (LSTM[29] and Gated Recurrent Unit (GRU[30] networks, which are specifically designed for sequential data processing.
 Shrestha et al.
@@
 Compared to CNN-based models in non-continuous tasks that utilize 135 M parameters [32] and RNN-based models in continuous tasks that utilize 189 M parameters [27] to achieve their best classification accuracy, hybrid models are more lightweight.
-However, the Floating-point Operation (FLOP) required by the hybrid model [14] is evaluated to be at least 1 G.
+However, the #\#FLOP/Inf.
+required by the hybrid model [17] is evaluated to be at least 1 G.
 The arithmetic intensity of CNNs, especially for multi-channel signals, drives their computational complexity.
@@
 Recently, researchers have explored transformer-based architectures and SSM in RadHAR.
-ViT-based methods [16] achieved 92% accuracy with 2.7 M parameters, while [21] reached 92.1% with 769 k parameters on non-continuous FMCW radar data [30], though both demand significant ##\##FLOP per inference.
+ViT-based methods [19] achieved 92% accuracy with 2.7 M parameters, while [33] reached 92.1% with 769 k parameters on non-continuous FMCW radar data [25], though both demand significant #\#FLOP per inference.
 (169 M and 2.41 G, respectively).
-In continuous HAR, however, ViT appears to under-perform conventional CNN-based models, such as the first to the third blocks of AlexNet and ResNet [17].
-SSM-based approaches [19] hit 99.82% accuracy on a non-continuous Continuous Wave (CW) dataset [30] with 8.7 M parameters and 1.22G ##\##FLOP per inference.
-III Analysis of Radar Data fed to ViM and Design Motivation
-Figure 1: The schematic of conventional Vision Mamba and its three potential issues while classifying the micro-Doppler-based human activity recognition.
-To effectively combine the modeling strength of SSM and decrease computational complexity, it is crucial to consider the unique properties of micro-Doppler signatures.
-To illustrate this, we trace the data flow from input to SSM in Vision Mamba (ViM), highlighting issues specific to radar micro-Doppler-based HAR tasks.
-Other details will be explained in section IV.
+In continuous HAR, however, ViT appears to underperform conventional CNN-based models, such as the first three blocks of AlexNet and ResNet [20].
+SSM-based approaches [22] hit 99.82% accuracy on a non-continuous CW dataset [25] with 8.7 M parameters and 1.22G #\#FLOP per inference.
+Compared to ultra-lightweight object detection models optimized for edge devices, such as YOLO-Fastest with 350 k parameters, and 252 M #FLOP/Inf, directly applying these advanced architectures to radar-based HAR tasks remains costly due to their high computational overhead.
+III Proposed RadMamba
+To effectively combine the modeling power of SSM with lower computational cost, it is essential to exploit the distinctive characteristics of micro-Doppler signatures.
+However, applying a conventional Vision Mamba (ViM) to RadHAR reveals three key issues:
+  • Issue 1: Micro-Doppler signatures exhibit significant background sparsity, so most time–frequency bins carry little task-relevant information.
+  • Issue 2: Rectangular patch segmentation disrupts the continuous spatio-temporal structure of micro-Doppler patterns, degrading representation quality.
+  • Issue 3: Standard linear projection layers struggle to capture Doppler-specific, time–frequency patterns.
+To clarify these points, we trace the data flow from input to SSM in RadMamba, highlighting issues specific to radar micro-Doppler-based HAR tasks and describing our proposed solutions in RadMamba.
 As shown in Fig.
-1, classical ViM comprises three core components: patch and position embedding, Mamba block with bidirectional branches, and SSMs.
-III-A Issue 1: Micro-Doppler signatures exhibit significant sparsity.
+1, RadMamba consists of three core components: preprocessing and segmentation; patch and positional embedding; and a convolution-projection–based Mamba block with bidirectional SSMs.
+We customize the architecture for micro-Doppler signals using three solutions: channel fusion with downsampling, Doppler-aligned segmentation, and the Convolutional Projection-based Mamba (CP-Mamba) block.
+Figure 1: The schematic of the proposed RadMamba architecture and its novel Solution 1 for radar micro-Doppler-based classification.
+III-A Solution 1: Channel Fusion and Downsampling
 The micro-Doppler signature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ serves as input to the classifier, with the Doppler frequency dimension represented as the image height $H$ and time as the image width $W$.
-The number of channels $C$ may correspond to channels from multiple radar sensors or one of the RGB channels.
-Analysis of Dataset [31] CI4R reveals that $\mathbf{X}$ exhibits an average spectrogram sparsity of approximately 87%, as illustrated in Fig.1 Issue 1.
-In the subsequent Segment step, $\mathbf{X}$ is divided into a matrix of smaller patches $\mathbf{X}_{seg}\in\mathbb{R}^{C\times I\times J\times H_{seg}\times W_{seg}}$, where $H_{seg}$ and $W_{seg}$ denote the height and width of each patch, respectively.
-Consequently, the number of patches per column is $I=\frac{H}{H_{seg}}$, and per row is $J=\frac{W}{W_{seg}}$.
-As depicted in Fig.1 Issue 1, many patches in $\mathbf{X}_{seg}$ contain only background information and resemble one another, providing little discriminative detail regarding Doppler frequency and time within each patch.
-III-B Issue 2: Destruction of Spatial and Temporal Information.
-Changes in Doppler frequency over time are critical for distinguishing various human activities.
-However, rectangular patch segmentation disrupts the relative positioning of Doppler frequency and time through two processes:
-III-B1 Segment Undermines Temporal Interpretability in SSM, as Shown in Fig. 1, Issue 2 Left
-As illustrated in Fig.1 left, a patch with column index $j$ can be interpreted as time point $t_{j}$, where $j=0,1,\cdots J$.
-During segmentation, the patch matrix of dimensions $I\times J$ is reshaped into a one-dimensional sequence $N=I\cdot J$.
-This reshaping results in a time point sequence of dimension $N$ as $[t_{0},t_{1},\cdots,t_{J},t_{0},t_{1},\cdots,t_{J},\cdots]$, represented by the yellow folded line in Fig.1 left.
-However, in a real-time sequence, $t_{J}$ and $t_{0}$ are not continuous.
-When the dimension $N$ remains unchanged from Segment to SSMbranch, and the SSM branches treat $N$ as the direction of temporal evolution, the SSM encounters a discontinuous sequence.
-This discontinuity hinders its ability to discern significant features in the micro-Doppler signature.
-III-B2 Flatten Compromises Spatial Information Integrity Across Time and Doppler Dimensions in Patches, as Shown in Fig. 1, Issue 2 Right
-Following segmentation, each patch is flattened into a vector:
-$\mathbf{X^{\prime}_{seg}}=\textit{Flatten}(\mathbf{X_{seg}})$ (1)
-where $\mathbf{X^{\prime}{seg}}\in\mathbb{R}^{N\times CH{seg}W_{seg}}$.
-A pixel with row index $h_{seg}$ can be interpreted as Doppler frequency point $f_{h_{seg}}$, where $h_{seg}=0,1,\cdots H_{seg}$.
-The Flatten operation results in a Doppler point sequence in $H_{seg}W_{seg}$ as $[f_{0},f_{1},\cdots,f_{h_{seg}},f_{0},f_{1},\cdots,f_{h_{seg}},\cdots]$, as shown in Fig.
-1 left.
-However, $f_{h_{seg}}$ and $f_{0}$ are not continuous in the Doppler frequency sequence, disrupting the Doppler frequency features.
-This disruption also affects the time dimension within each patch.
-In summary, the second issue encountered by conventional ViM is that Doppler and temporal information within $H_{seg}$ and $W_{seg}$ is compromised by flattening.
-III-C Issue 3: Challenges for Projection Layers in Capturing Doppler Patterns
-Within the Mamba block, as shown in Fig.
-1 top, the input $\mathbf{X_{p}}$ first undergoes layer normalization to yield $\mathbf{X}_{proj}$, which is then split into two branches via linear projections P1 and P2:
-$\mathbf{Z}$ $=\mathbf{W}_{\textit{P1}}\mathbf{X}_{proj}+\mathbf{b}_{\textit{P1}},$ (2)
-$\mathbf{X}_{fw}$ $=\mathbf{W}_{\textit{P2}}\mathbf{X}_{proj}+\mathbf{b}_{\textit{P2}},$ (3)
-where the $\mathbf{W}$ and $\mathbf{b}$ terms are the weight matrices and bias vectors, respectively.
-The projection operations are expected to extract Doppler frequency patterns over time and enhance the average cross-correlation between patches.
-This enhancement occurs because a high-performing classifier should recognize all Doppler-aligned vectors of a specific activity as a cohesive and highly relevant unit.
-However, the averaged cross-correlation between patches shows limited improvement from conventional projections P1 and P2.
-For instance, in Dataset CI4R [31], the average patch cross-correlation is 0.008 before projections 1 and 2.
-After projections 1 and 2, they rise to 77.9 and 0.4, respectively.
-In contrast to the improvements achieved through our proposed Conv1D projection (detailed later in Section IV-C), the linear layers in the conventional ViM architecture demonstrate limited capability to enhance cross-correlation between patches and consistently identify all time bins as components of the same action.
-The tensors $\mathbf{X}{fw}$ and $\mathbf{X}{bw}$ resulting from the projections P1 and P2 are then passed through SSM branches.
-The outputs from both directions are combined using a gating mechanism modulated by the SiLU-activation to produce the gated output $\mathbf{Y}{sum}$.
-This gated output passes through projection layer P3, which is linear in the conventional ViM, and is added to the input via a residual connection:
-$\mathbf{Y}_{out}=\mathbf{W}_{\textit{P3}}\mathbf{Y}_{sum}+\mathbf{b}_{\textit{P3}}+\mathbf{X}_{proj},$ (4)
-Before and after projection 3, again using Dataset CI4R as an example, the average patch cross-correlation is 160.1 and 75.5, respectively.
-Essentially, the averaged cross-correlation between patches is even decreased by the conventional projection layer 3, posing a challenge to achieve good classification performance.
-In the analysis of Issue 3, to compute the averaged patch cross-correlation, we first trained a model to its optimal performance, then duplicated and froze its parameters.
-By re-feeding the data into the frozen model, we calculated the averaged patch cross-correlation during inference.
-This averaged patch cross-correlation also accounts for different batches (classes).
-For example, with $\mathbf{X}{proj}\in\mathbb{R}^{B\times N\times D}$, each patch is labeled with its index $n$ in $\mathbf{X}{seg}$.
-The averaged patch cross-correlation is defined as:
+The channel dimension $C$ corresponds either to multiple radar signal channels or to RGB channels when visualized as an image.
+In the subsequent Segment step, $\mathbf{X}$ is partitioned into a matrix of smaller patches, denoted as $\mathbf{X}_{seg}$.
+As depicted in Fig.1 VisionMamba Issue 1, many patches in $\mathbf{X}_{seg}$ contain only background information and therefore appear visually redundant.
+Based on our analysis of the CI4R Dataset [26], the micro-Doppler spectrograms exhibit an average spectrogram sparsity of approximately 87%.
+Such sparse background regions contribute limited discriminative information regarding Doppler frequency or temporal dynamics, reducing the representational effectiveness of these patches.
+To address the high sparsity inherent in micro-Doppler signatures, we introduce a preprocessing pipeline, Channel Fusion and Downsampling (Chan-DS), before the segmentation stage, as depicted in Fig.
+1.
+$\mathbf{X_{c-d}=\text{Chan-DS}(\mathbf{X})}$ (1)
+where $\mathbf{X}_{c-d}\in\mathbb{R}^{C_{c-d}\times H_{c-d}\times W_{c-d}}$ with $H_{c-d}<H$ and $W_{c-d}<W$ is the micro-Doppler signature after channel fusion and downsampling.
+The architecture of the Chan-DS block is as follows:
+$\text{Chan-DS}=[$ $(\text{Conv2D},\text{BatchNorm})\times L,$
+$\text{Maxpool2D},$
+$\text{Maxpool1D},$
+$\text{Avgpool1D}\text{ (optional)}]$
+where $L=1,2$.
+The Two-Dimensional (2D) convolutional layer and batch normalization layer mix channel information and extract preliminary features from the input tensor $\mathbf{X}$.
+A 2D max pooling layer and a One-Dimensional (1D) max pooling layer reduce spatial dimensions.
+Optionally, an additional 1D average pooling layer further downsamples the time dimension if the data sparsity along the time axis is excessive for FMCW radar datasets.
+This approach avoids generating numerous empty or redundant patches.
+Additionally, downsampling reduces the spatial dimensions before segmentation, thereby further contributing to an energy-efficient design.
+III-B Solution 2: Doppler-Aligned Segmentation and Patch Embedding
+Figure 2: VisionMamba Issue 2 and RadMamba Solution 2: Doppler-Aligned Segmentation
+The feature map after segmentation $\mathbf{X}_{seg}$ has dimensions of $C\times I\times J\times H_{seg}\times W_{seg}$, where $H_{seg}$ and $W_{seg}$ denote the height and width of each patch.
+Thus, the number of patches per column and per row are determined by $I=\frac{H}{H_{seg}}$ and $J=\frac{W}{W_{seg}}$, respectively, resulting in a total number of patches $N=IJ$.
+As illustrated in Fig.
+2 (a) (upper), patches indexed by $j$ along the horizontal direction correspond to temporal positions $t_{j}$, where $j=0,1,\ldots,J$.
+Within each patch, as shown in Fig.
+2 (a) (bottom), pixels indexed by $h_{seg}$ along the vertical direction correspond to Doppler frequency points $f_{h_{seg}}$, where $h_{seg}=0,1,\ldots,H_{seg}$.
+Each patch is then flattened into a vector:
+$\mathbf{X^{\prime}_{seg}}=\text{Flatten}(\mathbf{X_{seg}})$ (2)
+where $\mathbf{X^{\prime}_{seg}}\in\mathbb{R}^{N\times CH_{seg}W_{seg}}$.
+The second limitation of conventional VisionMamba lies in the fact that patch flattening disrupts Doppler–temporal continuity, as depicted in Fig.
+2 (a).
+On the one hand, rectangular segmentation results in a temporal ordering in $\mathbf{X}^{\prime}_{seg}$ along the dimension $N$ of the form:
+$[t_{0},t_{1},\cdots,t_{J},t_{0},t_{1},\cdots,t_{J},\cdots]$ (3)
+represented by the yellow folded arrow in Fig.
+2 (a).
+Since VisionMamba treats $N$ as a temporal evolution dimension in the SSM branches, the artificial jump from $t_{J}$ to $t_{0}$ introduces a temporal discontinuity, making it difficult for the SSM to extract meaningful temporal dynamics.
+On the other hand, the flattening operation yields a Doppler frequency sequence inside each patch of the form:
+$[f_{0},f_{1},\cdots,f_{h_{seg}},f_{0},f_{1},\cdots,f_{h_{seg}},\cdots]$ (4)
+as shown in Fig.
+2 bottom.
+The discontinuity between $f_{h_{seg}}$ and $f_{0}$ breaks the frequency pattern within each patch, simultaneously degrading both Doppler frequency consistency and the temporal evolution represented in that patch.
+To address this issue, we propose Doppler-aligned segmentation, as shown in Fig. 2 (b):
+$\mathbf{X_{seg}}=\text{DA-Seg}(\mathbf{X}_{c-d})$ (5)
+Specifically, we define patches with a height of $H_{seg}=H_{c-d}$ (the full Doppler dimension) and a width of $W_{seg}=1$ (a single time bin of the spectrogram).
+This Doppler-aligned segmentation ensures that the temporal evolution of the patch sequence remains strictly ordered as:
+$[t_{0},t_{1},\cdots,t_{J}]$ (6)
+corresponding directly to the time sequence $W$.
+When the segmented data is fed into the SSM, as shown in Fig.
+2 (b) (upper), the SSM receives a continuous temporal input stream.
+Furthermore, Doppler-aligned segmentation produces a patch vector with dimension
+$CH_{seg}W_{seg}=H_{c-d}$ (7)
+preserving the entire Doppler profile within each patch. Consequently, after flattening, the Doppler frequency sequence remains intact as:
+$[f_{0},f_{1},\cdots,f_{h\_seg}]$ (8)
+as illustrated in Fig. 2(b) (bottom). This maintains Doppler continuity, enabling the model to leverage fine-grained frequency dynamics without disruption.
+III-C Solution 3: Convolution Projections in CP-Mamba Blocks
+Within the conventional Mamba block (Fig. 1), the input $\mathbf{X}_{p}$ is first normalized to obtain $\mathbf{X}_{proj}$, which is then split into three paths:
+$\mathbf{Z}$ $=\text{Projection}_{P1}(\mathbf{X}_{proj}),$ (9)
+$\mathbf{X}_{fw}$ $=\text{Projection}_{P2}(\mathbf{X}_{proj}),$ (10)
+$\mathbf{X}_{bw}$ $=\text{Flip}(\mathbf{X}_{fw},dim=\text{temporal dim.}),$ (11)
+where $\mathbf{Z},\mathbf{X}_{fw},\mathbf{X}_{bw}\in\mathbb{R}^{N\times dim}$, and Flip() reverses the temporal dimension of $\mathbf{X}_{fw}$.
+In the conventional Mamba block, $\text{Projection}_{P1}$ and $\text{Projection}_{P2}$ are linear layers.
+In contrast, the CP-Mamba block replaces them with 1D convolutions using kernel size $(dim,dim,3)$.
+Next, $\mathbf{X}_{fw}$ and $\mathbf{X}_{bw}$ are processed by the two SSM branches:
+$[\text{Conv1D}{fw/bw},\text{LayerNorm}{fw/bw},\text{SSM}{fw/bw}],$ (12)
+where $\text{Conv1D}{fw/bw}$ has weight shape $(dim,dim,1)$ and operates along the sequence dimension $(N,dim)$.
+The two directional outputs are gated by the Sigmoid Linear Unit (SiLU)-activated $\mathbf{Z}$ and summed:
+$\mathbf{Y}^{\prime}{fw/bw}$ $=\mathbf{Y}{fw/bw}\odot\text{SiLU}(\mathbf{Z}),$ (13)
+$\mathbf{Y}{sum}$ $=\mathbf{Y}^{\prime}{fw}+\mathbf{Y}^{\prime}{bw},$ (14)
+where $\odot$ denotes element-wise multiplication. Finally, a projection P3 and residual connection produce the output:
+$\mathbf{Y}_{out}=\text{Projection}_{P3}(\mathbf{Y}_{sum})+\mathbf{X}_{proj},$ (15)
+In the conventional Mamba block, the projection layer P3 is linear.
+In our CP-Mamba, P3 is replaced with a Conv1D, where $\text{Conv1D}_{P3}$ has weights $(dim,dim,1)$ and restores the sequence shape to $(N,dim)$.
+TABLE I: Averaged cross-correlation between patches before and after each projection layer.
+Linear Projection Layers
+in Conventional Mamba
+Convolutional Projection Layers
+in Our Proposed CP-Mamba
+Projection At Input At Output At Input At Output
+1 0.008 77.9 0.002 1065.7
+2 0.008 0.4 0.002 16.1
+3 160.1 75.5 245.3 356.4
+Ideally, projection layers should improve cross-correlation between patches, since a strong classifier should treat Doppler-aligned patches of the same activity as a coherent unit.
+Linear projections, however, increase correlation only marginally (P1, P2) or even reduce it (P3), limiting classification performance.
+The convolutional projections in CP-Mamba consistently strengthen cross-correlation, making patches easier to identify as belonging to the same action and improving Doppler-pattern sensitivity.
+Tab.
+I reports the averaged patch cross-correlation using the Dataset CI4R [26].
+For conventional linear projections, the averaged correlation before projections 1 and 2 is 0.008, rising to 0.77 and 0.4, respectively.
+Before and after projection 3, it drops from 160.1 to 75.5.
+In contrast, our convolutional projections greatly enlarge the correlation gap (e.g., from 77.9 to 1065.7 for P1), indicating superior ability to aggregate time bins of the same activity.
+To compute the averaged patch cross-correlation, we train the model, freeze the parameters, and refeed the data.
+For $\mathbf{X}_{proj}\in\mathbb{R}^{B\times N\times D}$, each patch is indexed by $n$.
+The averaged cross-correlation is defined as:
 $Corr_{avg}=$ $\frac{1}{B}(\frac{1}{N}\sum_{n=1}^{N}(\frac{1}{N}\sum_{n^{\prime}=1}^{N}($
@@
 $corr(\mathbf{A},\mathbf{B})=\sum_{m=0}^{M}\sum_{m^{\prime}=0}^{M}\mathbf{A}[m]\cdot\mathbf{B}[m+m^{\prime}]$ (17)
-IV Proposed RadMamba
-To overcome the aforementioned challenges and issues, we propose a micro-Doppler-oriented RadMamba architecture, illustrated in Fig. 2.
-Figure 2: The schematic of the proposed RadMamba architecture and its novel solutions for radar micro-Doppler-based classification.
-These modifications to the conventional approach include channel fusion and down-sampling, Doppler-aligned segmentation, and modified Mamba block rMamba.
-IV-A Solution 1: Channel Fusion and Downsample
-To tackle the issue of many similar and empty patches caused by high sparsity inherent in micro-Doppler signatures input $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, we introduce a preprocessing pipeline prior to the segmentation stage, as depicted in Fig.
-2 Chan-DS: Channel Fusion and Downsample.
-$\mathbf{X_{c-d}=\textit{Chan-DS}(\mathbf{X})}$ (7)
-where $\mathbf{X}\in\mathbb{R}^{C_{c-d}\times H_{c-d}\times W_{c-d}},H_{c-d}<H,W_{c-d}<W$ is the input micro-Doppler signature after the channel fusion and downsample for each batch.
-The architecture of the Chan-DS block is as follows:
-[…]
-[…]
-[…]
-[…]
-where $L=1,2$.
-The Two-Dimensional (2D) convolutional layer and batch normalization layer mix channel information and extract preliminary features from the input tensor $\mathbf{X}$.
-A 2D max pooling layer and a One-Dimensional (1D) max pooling layer reduces spatial dimensions.
-Optionally, an additional 1D average pooling layer with a kernel height as one and kernel width larger than 1, for example (1, 4), further downsamples the time dimension if the data sparsity along the time axis, influenced by the Pulse Repetition Interval (PRI) of the radar and pre-processing to generate the input, is excessive.
-In combination with the Doppler-Aligned Segment explained in the next subsection IV-B, this method avoids the issue of having many similar and empty patches of data.
-Moreover, the down-sampling decreases the available size for patch segmentation, which enhances the energy-efficient design.
-IV-B Solution 2: Doppler-Aligned Segment and Patch Embedding
-To maintain the temporal and Doppler location information embedded in micro-Doppler signatures, we propose a new patch segmentation strategy, as shown in Fig.
-2 Solution 2 Doppler-Aligned Segment, DA-Seg:
-$\mathbf{X_{seg}}=\textit{DA-Seg}(\mathbf{X}_{c-d})$ (8)
-Specifically, we define patches with a height of $H_{seg}=H_{c-d}$ (the full Doppler dimension) and a width of $W_{seg}=1$ (a single time bin of the spectrogram).
-This approach ensures that the temporal evolution of state sequence $N$ remains as $[t_{0},t_{1},\cdots,t_{J}]$ and continuous as time sequence $W$.
-When the data is sent to the SSM, as depicted in Fig.2 Solution 2 left, this method avoids the loss of critical temporal information that conventional square patches of the ViM might face.
-Moreover, this segmentation approach results in a pixel sequence dimension $CH_{seg}W_{seg}=H_{c-d}$ of $\mathbf{X}^{\prime}_{seg}$ that preserves the frequency pattern as $[f_{0},f_{1},\cdots,f_{h\_seg}]$, during Flatten operations within each Doppler bin, as shown in Fig.2 Solution 2 left, as:
-$\mathbf{X^{\prime}_{seg}}=\textit{Flatten}(\mathbf{X_{seg}})=\mathbf{X_{seg}}$ (9)
-IV-C Solution 3: Convolution Projections in rMamba Blocks
-Finally, to increase the capability of projection layers to capture the Doppler patterns, we replace the conventional linear projection layers within the conventional Mamba block with 1D convolutional layers, as highlighted in Fig.
-2 rMamba Block, in the blue boxes on the top-right part of the figure.
-Within the rMamba block, the linear projections P1 and P2 are replaced by Conv1D layer as:
-[…]
-[…]
-$\mathbf{X}_{bw}$ $=\text{flip state dim}(\mathbf{X}_{fw}),$ (12)
-where $\mathbf{Z},\mathbf{X}_{fw},\mathbf{X}_{bw}\in\mathbb{R}^{N\times dim}$. Both Conv1D operations utilize a kernel size (in channels, out channels, kernel size) = $(dim,dim,3)$.
-When $\mathbf{X}_{fw}$ and $\mathbf{X}_{bw}$ pass through the SSM branches with architecture of:
-[…]
-[…]
-The outputs from both directions are combined using a gating mechanism modulated by the SiLU-activated $\mathbf{Z}$ and summed:
-$\mathbf{Y^{\prime}}_{fw/bw}$ $=\mathbf{Y}_{fw/bw}\odot\text{SiLU}(\mathbf{Z})$ (14)
-$\mathbf{Y}_{sum}$ $=\mathbf{Y^{\prime}}_{fw}+\mathbf{Y^{\prime}}_{bw}$ (15)
-where $\odot$ denotes element-wise multiplication. Additionally, in our proposed rMamba, the projection layer P3 is also replaced by Conv1D, so that:
-$\mathbf{Y}_{out}=\text{Conv1D}_{P3}(\mathbf{Y}_{sum})+\mathbf{X}_{proj},$ (16)
-where the final Conv1D has weights of dimensions $(dim,dim,1)$ and transforms the combined sequence back to $(N,dim)$.
-These convolutional layers enable the model to capture subtle variations in Doppler-span changes over time, which are crucial to describe the kinematic evolution of the different activities to be classified.
-Also, let us take Dataset CI4R as a comparison reference and follow the method to calculate cross-correlation as explained in Section III Issue 3.
-While passing through RadMamba, the average time bin cross-correlation is 0.002 before projection 1 and 2.
-Then, compared to conventional linear projection, the convolution projection 1 increased the averaged cross correlation gap between projection input and output from 77.9 to 1065.7.
-The convolutional projection 2 increased this gap from 0.4 to 16.1.
-The convolutional projection 3 increased the average time cross-correlation from 245.3 to 356.4, while the linear projection 3 decreased this number.
-This enhancement significantly improves the model’s sensitivity to Doppler patterns between patches and makes it easier to regard all patches as one activity, thereby boosting classification performance.
-IV-D Selective State-Space Modeling for Micro-Doppler Sequences
-Figure 3: The architecture of State Space Model in RadMamba.
-The SSM is a significant layer in rMamba enables efficient sequence modeling for the dynamic nature of micro-Doppler signatures. As shown in Fig. 3 (a), the discrete SSM is defined as:
+III-D Selective State-Space Modeling for Micro-Doppler Sequences
+Figure 3: The architecture of the State Space Model in RadMamba.
+The SSM is a significant layer in CP-Mamba that enables efficient sequence modeling for the dynamic nature of micro-Doppler signatures. As shown in Fig. 3 (a), the discrete SSM is defined as:
 $\mathbf{h}[n]$ $={\mathbf{\bar{A}}}[n]\mathbf{h}[n-1]+{\mathbf{\bar{B}}}[n]\mathbf{x}[n],$ (18)
@@
 $\bar{\mathbf{A}}$ $=\exp(\boldsymbol{\Delta}\cdot\mathbf{A}),$ (20)
-[…]
-The input of the SSM branches in the proposed rMamba block are tensors $\mathbf{X}_{fw/bw}\in\mathbb{R}^{N\times dim}$; after Conv1D and LayerNorm in SSM branch, this is termed as $\mathbf{X^{\prime}}_{fw/bw}\in\mathbb{R}^{N\times dim}$.
-Therefore, the dimension of the discretized state transition ${\mathbf{\bar{A}}}$ is ${N\times dim\times{dim_{s}}}$ and the input matrices dimension ${\mathbf{\bar{B}}}$ is ${N\times N\times{dim_{s}}}$, respectively.
-While $N$ equals the time dimension in micro-Doppler signatures, the input matrices $\mathbf{\bar{B}}$ captures how the Doppler-aligned vector influences the state $\mathbf{h}[n]$, and the discretized state transition weight ${\mathbf{\bar{A}}}$ capture how the last state $\mathbf{h}[n-1]$.
+$\bar{\mathbf{B}}$ $=\left(\int_{0}^{\boldsymbol{\Delta}}\exp(\mathbf{A}\tau)\,d\tau\right)\mathbf{B},$ (21)
+The input of the SSM branches in the proposed CP-Mamba block are tensors $\mathbf{X}_{fw/bw}\in\mathbb{R}^{N\times dim}$; after Conv1D and LayerNorm in SSM branch, this is termed as $\mathbf{X^{\prime}}_{fw/bw}\in\mathbb{R}^{N\times dim}$.
+Therefore, the discretized state-transition tensor $\mathbf{\bar{A}}$} has dimension $N\times dim\times dim_{s}$, and the input tensor $\mathbf{\bar{B}}$} has dimension $N\times N\times dim_{s}$.
+Here, $N$ equals the time dimension in the micro-Doppler signatures.
+The input tensor $\mathbf{\bar{B}}$ captures how the Doppler-aligned input vector influences the state $\mathbf{h}[n]$, whereas the state-transition tensor $\mathbf{\bar{A}}$} captures how the previous state $\mathbf{h}[n-1]$ contributes to $\mathbf{h}[n]$.
 Moreover, the matrices $\mathbf{C}\in\mathbb{R}^{N\times dim_{s}}$ translate the current state $\mathbf{h}[n]$ to the output.
-The skip matrices $\mathbf{D}\in\mathbb{R}^{dim}$ add weighted input directly to output of SSM, which helps the recurrent part focus on the temporal evolution.
+The skip matrices $\mathbf{D}\in\mathbb{R}^{dim}$ add weighted input directly to the output of SSM, which helps the recurrent part focus on the temporal evolution.
 The selective mechanism in the SSM introduces input-dependent parameterization that enhances adaptability as shown in Fig.
@@
 $\boldsymbol{\Delta}$ $=\text{Softplus}(\mathbf{W}_{\boldsymbol{\Delta}}\mathbf{X}),$ (24)
-[…]
-IV-E Overall Architecture
+where $\mathbf{W}_{\mathbf{B}}$ and $\mathbf{W}_{\mathbf{C}}$ are weights of dimension $(dim,dim_{s})$.
+The $\boldsymbol{\Delta}$ computation involves a linear layer with weights $(dim,\text{dt\_rank})$ followed by a projection to $(N,dim)$ via weights $(\text{dt\_rank},dim)$, where dt_rank is a hyperparameter controlling the rank of the dynamic tensor.
+The Softplus activation ensures positive step sizes, implemented as $\text{Softplus}(x)=\log(1+\exp(x))$.
+III-E Overall Architecture
 To summarize the method, the proposed RadMamba pipeline in pseudo-code is presented in Algorithm I.
-The patch embedding in PEmb follows the method in  [15] and the position embedding method is a sinusoidal position encoding.
+The patch embedding in PEmb follows the method in  [18], and the position embedding method is a sinusoidal position encoding.
 [34].
@@
 Input: Micro-Doppler Signature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$
-Parameters: $dim$, $dim_{s}$, $dt\_rank$
+Parameters: dimdim, $dim_{s}$, dt_rankdt\_rank
 Output: Classification Result $\mathbf{F}\in\mathbb{R}^{Q}$, where $Q$ is the number of classes
 1:Channel Fusion and Downsample:
-2:   $\mathbf{X_{c-d}}=\textit{Chan-DS}(\mathbf{X})$, $\mathbf{X}_{c-d}\in\mathbb{R}^{C_{c-d}\times H_{c-d}\times W_{c-d}}$
+[…]
 3:Doppler-Aligned Segmentation:
-4:   $\mathbf{X_{seg}}=\textit{DA-Seg}(\mathbf{X}_{c-d})$, $\mathbf{X}_{seg}\in\mathbb{R}^{C_{c-d}\times N\times H_{c-d}}$
+[…]
 5:Patch and Position Embedding:
-6:   $\mathbf{X^{\prime}}_{p}=\textit{PEmb}(\mathbf{X}_{seg})$, $\mathbf{X^{\prime}}_{p}\in\mathbb{R}^{N\times D}$
-7:rMamba Block:
-8:   $\mathbf{Y}_{out}=\textit{rMamba}(\mathbf{X}_{p})$, $\mathbf{Y}_{out}\in\mathbb{R}^{N\times dim}$
+[…]
+7:CP-Mamba Block:
+[…]
 9:Classify:
-10:   $\mathbf{F}=\mathbf{W}_{out}\mathbf{Y}_{out}$, $\mathbf{F}\in\mathbb{R}^{Q}$
-V Open-source Datasets and Preprocessing for Radar-based HAR
+[…]
+IV Open-source Datasets and Preprocessing for Radar-based HAR
 To evaluate the efficacy of our proposed RadMamba framework, we utilize three distinct radar datasets.
-This section describes these datasets, namely the CW non-continuous Dataset DIAT, FMCW non-continuous Dataset CI4R, and FMCW continuous Dataset UoG2020, along with their respective pre-processing steps.
+This section describes these datasets, namely the CW non-continuous Dataset DIAT, FMCW non-continuous Dataset CI4R, and FMCW continuous Dataset UoG20, along with their respective pre-processing steps.
 More details on these datasets are summarized in Tab.
-I.
-Here the word ‘continuous’ refers to samples of data where sequences of human activities are performed one after the other in a continuous manner.
-TABLE I: Information of Different Radar-based Activity Recognition Datasets Used in This Work.
-Datasets DIAT CI4R UoG2020
+II.
+Here, the word ‘continuous’ refers to samples of data where sequences of human activities are continuously performed one after the other.
+Examples of micro-doppler signature after pre-processing are shown in Fig.
+4.
+TABLE II: Overview of radar-based human activity recognition datasets used in this work.
+Datasets DIAT [25] CI4R [26] UoG20 [27]
 Continuous No No Yes
@@
 Freq.(GHz) 10 77 5.8
-Resolution 224$\times$224 224$\times$224 240$\times$1471
-BW(MHz) - 750 400
+[…]
+BW(MHz) N/A 750 400
 Length per sample (s) 3 20 35
@@
 Test Samples 759 132 6
-Sliding Window 224
-V-A Dataset DIAT: X-Band CW Radar with Discrete Activity Instances [30]
+Sliding Window N/A N/A 224
+Figure 4: Micro-Doppler Signature Example of (a) Dataset DIAT (b) Dataset CI4R (c) Dataset UoG20
+IV-A Dataset DIAT: X-Band CW Radar with Discrete Activity Instances [25]
 This dataset was collected using an X-band CW radar operating at a frequency of 10 GHz.
-It encompasses data from 30 human subjects performing ‘suspicious’ activities at ranges spanning 10 meters to $\pm$0.5 kilometers over 3-second durations, resulting in an imbalanced set of 3,780 spectrogram samples of shape $(3,244,244)$across six activity classes.
+It encompasses data from 30 human subjects performing ‘suspicious’ activities at ranges spanning 10 meters to $\pm$0.5 kilometers over 3-second durations, resulting in an imbalanced set of 3,780 spectrogram samples of shape $(3,224,224)$ across six activity classes.
 Preprocessing adheres to the methodology outlined in [25].
 To facilitate a fair comparison with previous works, the dataset is partitioned into 70% training, 10% validation, and 20% test sets within each class, maintaining class-specific balance for reliable performance evaluation.
-V-B Dataset CI4R: Short-Range FMCW Radar for Multi-Class Activity Recognition [31]
-The experiments of the second dataset employ a Texas Instruments IWR1443 FMCW radar, operating at 77 GHz with a bandwidth of 750 MHz.
+IV-B Dataset CI4R: Short-Range FMCW Radar for Multi-Class Activity Recognition [26]
+For the second dataset, experiments employ a Texas Instruments IWR1443 FMCW radar, operating at 77 GHz with a bandwidth of 750 MHz.
 It includes data from six participants, varying in age, height, and weight, performing 11 distinct activities and ambulatory gaits.
 Each participant executed 10 repetitions per activity, yielding 60 samples per class per sensor, as detailed in [26].
-Preprocessing involves applying a 256-point short-time Fourier transform (STFT) with a context window length to convert the raw radar signals into a 2D micro-Doppler representation of shape $(1,244,244)$.
+Preprocessing involves applying a 256-point short-time Fourier transform ( Short-Time Fourier Transform (STFT)) with a context window length of 256 to convert the raw radar signals into a 2D micro-Doppler representation of shape $(1,224,224)$.
 The dataset is split with an 8:2 training-to-test ratio, ensuring sufficient data for model optimization and evaluation.
-V-C Dataset UoG2020: Continuous FMCW Radar for Realistic HAR Sequences [27]
+IV-C Dataset UoG2020: Continuous FMCW Radar for Realistic HAR Sequences [27]
 Representing a leap in complexity, this dataset was acquired using an FMCW radar operating at 5.8 GHz with a 400 MHz bandwidth.
@@
 These activities are organized into three distinct sequential orders, with participants granted autonomy in selecting transition points between consecutive activities and determining trajectories for translational tasks, as described in [27].
-The resulting micro-Doppler signatures are preprocessed into $(1,240,244)$ 2D frames, with two participants’ data designated as the test set, while the remaining 13 contribute to training, providing a challenging benchmark for continuous sequence modeling.
-VI Experimental Setup
-VI-A Training Hyperparameter
-Training is conducted using the ADAMW optimizer [33, 34], paired with the ReduceLROnPlateau scheduler to adjust the learning rate based on test performance adaptively.
+The resulting micro-Doppler signatures are preprocessed into $(1,240,224)$ 2D frames, with data from two participants designated as the test set and data from the remaining 13 participants used for training, providing a challenging benchmark for continuous sequence modeling.
+V Experimental Setup
+V-A Network Structure
+TABLE III: Network Architecture of Models Used in This Work
+Model Swept Hidden Size (dimdim for RadMamba)
+ResNet6 (Customized) 5, 7, 12, 20, 30
+Bi-LSTM 1, 2, 4, 10, 20, 28, 34
+CNN-LSTM 2, 4, 6, 16, 28, 34, 46
+CNN-Bi-GRU 1, 2, 4, 6, 16, 28
+RadMamba Dataset DIAT 8, 16, 32, 64, 80
+RadMamba Dataset CI4R 8, 16, 32, 64, 80, 96, 128, 160
+RadMamba Dataset UoG20 8, 16, 20, 24, 32
+The classification performance of RadMamba is evaluated against a suite of established models, selected for their proven efficacy across CNN, RNN, and hybrid architectures: customized ResNet6 [35], Bi-LSTM [27], CNN-LSTM [16], and CNN-Bi-GRU [17].
+To explore the impact of model scaling, we conducted a model size sweep experiment, varying the hidden dimensions as shown in Tab.
+III.
+These sizes are chosen to capture the full spectrum of performance trends as model capacity increases, providing insights into scalability and efficiency.
+The customized ResNet6 is derived from ResNet18 by preserving the first five layers and the final classifier while maintaining its channel ratio between consecutive channels.
+The minimum channel width is tunable to target low-FLOP regimes, enabling fair comparisons under limited computation conditions.
+Although the CNN-Bi-GRU retains a larger parameter budget, its fusion strategy may not generalize consistently across datasets.
+All four baselines (ResNet6, Bi-LSTM, CNN-LSTM, and CNN-Bi-GRU) are retrained from scratch in a unified PyTorch pipeline, using consistent dataset splits, preprocessing, optimization settings, and batch=1 memory profiling.
+For transformer and SSM competitors, FML-ViT [19], LW-ViT [33], and ActivityMamba [22], we report their best published small-scale configurations.
+Following their protocols, we adopt the same training settings for FML-ViT and ActivityMamba, but use a different input size than LW-ViT: their preprocessing produces 656×656 inputs on the CI4R dataset, whereas our pipeline uses 224×224.
+TABLE IV: Specified Network Architecture Variations of RadMamba
+Datasets $dim_{s}$ dt_rankdt\_rank Chan. Fusion DS Redu. Factor
+Dataset DIAT 1 2 2, (16, 3, 3) (2, 2)
+Dataset CI4R 4 0 1, (1, 3, 3) (2, 8)
+Dataset UoG20 16 4 1, (1, 3, 3) (2, 32)
+For RadMamba, the configurations of channel fusion, downsampling, and SSM parameters are dataset-specific to optimize micro-Doppler processing (Tab.
+IV).
+The ‘channel fusion’ column specifies [L, ($C_{c-d},H_{k},W_{k}$)], where L is the number of channel fusion blocks and the $C_{c-d},H_{k},W_{k}$ are channel, kernel height, kernel width of Conv2D layer, respectively.
+The ‘DS Redu.
+Factor’ denotes the reduction factors ($H/H_{c-d}$, $W/W_{c-d}$).
+Hyperparameter selection follows two principles:
+  • Bayesian optimization: Because RadMamba is highly configurable, we perform a Bayesian-optimization-based search over a constrained space.
+Since dimdim predominantly determines the sizes of most weight matrices, the number of channels in Chan-DS, $dim_{s}$, and $dt_{rank}$ were selected to minimize model complexity across different dimdim.
+  • Physical perspective from micro-Doppler signature (Temporal downsampling and Reduction Factor): The first downsampling factor of $2\times 2$ in Chan-DS matches the fact that most Doppler energy appears in the central part of the spectrum.
+With horizontal reduction factors of $8$ and $32$ for the CI4R and UoG20 datasets (as shown in Tab.
+IV), each Doppler vector spans approximately 640  ms.
+Empirically, this temporal interval provides an optimal balance for both FMCW datasets.
+V-B Training Hyperparameter
+Training is conducted using the Adaptive Moment Estimation with decoupled Weight decay (AdamW) optimizer [36, 37], paired with the ReduceLROnPlateau scheduler, which adjusts the learning rate adaptively based on test performance.
 For Dataset DIAT, the initial learning rate is set to $1\times 10^{-4}$, with a batch size of 16, and training proceeds for 50 epochs.
 For Dataset CI4R, RadMamba employs an initial learning rate of $5\times 10^{-3}$, while the baseline models use $1\times 10^{-4}$, with a batch size of 16 over 100 epochs, reflecting the dataset’s increased complexity and sample size.
-For Dataset UoG2020, the initial learning rate is $5\times 10^{-5}$, with a larger batch size of 256 and 50 epochs, processing continuous sequences with a frame length of 224 and a stride of 1 to capture temporal dependencies effectively.
+For Dataset UoG20, the initial learning rate is $5\times 10^{-5}$, with a larger batch size of 256 and 50 epochs, processing continuous sequences with a frame length of 224 and a stride of 1 to effectively capture temporal dependencies.
 All experiments are executed on a single NVIDIA RTX 4090 GPU.
-To account for training variability, all models, including RadMamba, are trained and evaluated ten times with distinct seeds, reporting in the following analysis the mean and standard deviation of performance metrics in the model size swept experiment.
-Figure 4: Accuracy vs. number of parameters (1) and number of ##\##FLOP per inference (2) on: (a) Dataset DIAT (b) Dataset CI4R (c) Dataset UoG2020
-VI-B Network Structure
-TABLE II: Network Architecture of Models Used in This Work
-Model Swept Hidden Size ($dim$ for RadMamba)
-ResNet [5, 7, 12, 20, 30]
-Bi-LSTM [1, 2, 4, 10, 20, 28, 34]
-CNN-LSTM [2, 4, 6, 16, 28, 34, 46]
-CNN-Bi-GRU [1, 2, 4, 6, 16, 28]
-RadMamba Dataset DIAT [8, 16, 32, 64, 80]
-RadMamba Dataset CI4R [8, 16, 32, 64, 80, 96, 128, 160]
-RadMamba Dataset UoG2020 [8, 16, 20, 24, 32]
-TABLE III: Specified Network Architecture Variations of RadMamba
-Datasets $dim_{s}$ $dt\_rank$ Chan. Fusion DS Redu. Factor
-Dataset DIAT 1 2 [2, (16, 3, 3)] (2, 2)
-Dataset CI4R 4 0 [1, (1, 3, 3)] (2, 8)
-Dataset UoG2020 16 4 [1, (1, 3, 3)] (2, 32)
-The classification performance of RadMamba is evaluated against a suite of established models, selected for their proven efficacy across CNN, RNN, and hybrid architectures: ResNet [35], Bi-LSTM [27], CNN-LSTM [13], and CNN-Bi-GRU [14].
-To explore the impact of model scaling, we conducted a model size swept experiment, varying the hidden dimensions as shown in Tab.
-III.
-These sizes are chosen to capture the full spectrum of performance trends as model capacity increases, providing insights into scalability and efficiency.
-Notably, the CNN-Bi-GRU model, originally designed for distributed radar systems in  [14], exhibits a larger parameter count, though its channel fusion approach may not generalize across all datasets.
-For RadMamba, specific configurations of the channel fusion and down-sampling blocks, as well as the SSM parameters, are tailored to each dataset to optimize micro-Doppler processing as shown in Tab.
-III, where the ‘channel fusion’ column specifies [L, ($C_{c-d},H_{k},W_{k}$)], where L is the number of channel fusion blocks and the $C_{c-d},H_{k},W_{k}$ are channel, kernel height, kernel width of Conv2D layer, respectively.
-The ‘DS Redu.
-Factor’ denotes the reduction factors ($H/H_{c-d}$, $W/W_{c-d}$).
-For transformer-based and SSM-based models such as FML-ViT [16], LW-ViT [21], and ActivityMamba [19], we adopt their reported results as optimal outcomes under minimal model sizes, as dimension definitions may vary and are not always fully specified in the original works.
-VII Results and Discussion
-VII-A Model Scaling and Efficiency Analysis
-The model size swept results, illustrated in Figure 4, compare accuracy against the number of parameters from 2k to 150k and ##\##FLOP per inference across Datasets DIAT, CI4R, and UoG2020 between previous models ResNet, Bi-LSTM, CNN-Bi-GRU, and CNN-LSTM and the proposed RadMamba.
-For Dataset DIAT, as presented in Fig.
-4 (a1) and (a2), RadMamba achieves a peak accuracy of 99.8% with only 21.7 k parameters.
-However, RadMamba overfits with a model size larger than 21.7 k because the dataset is very simple and the accuracy is close to 100%.
-In contrast, ResNet’s accuracy increases steadily with model size, reaching 99.9% with 208.1 k parameters, yet it requires nearly ten times the parameters of RadMamba for slightly lower performance.
-TABLE IV: Mean and Standard Deviation of Classification Accuracy Performance across Seed 0 to 9 of Different NN-based RadHAR Models Evaluated With Three Different Radar Settings alongside Their Model Size and Floating-point Operations per Inference Sample (FLOP/Inf.)
+To account for training variability, all models, including RadMamba, are trained and evaluated ten times with distinct seeds.
+The following analysis reports the mean and standard deviation of performance metrics from the model size swept experiment.
+V-C Evaluation Settings
+Inference time and memory profiling are conducted on an NVIDIA Jetson Orin Nano Developer Kit with PyTorch 2.8.0 and CUDA 12.6.
+The kit is equipped with a 6-core Arm Cortex-A78AE CPU and an NVIDIA Ampere Architecture GPU with unified off-chip 8 GB 128-bit LPDDR5 memory.
+The evaluation follows a standardized protocol: each model undergoes 10 warm-up iterations followed by 100 measurement iterations, with inference performed at a batch size of 1.
+Peak memory (Peak Mem.) is measured using torch.cuda.memory_allocated(device), which returns the maximum allocated memory across all 100 inference iterations.
+CUDA synchronization with the ARM CPU is enforced before and after each forward pass to ensure accurate timing measurements.
+VI Results and Discussion
+VI-A Comparison with Previous Works
+TABLE V: Mean and Standard Deviation of Classification Accuracy Performance across Seed 0 to 9 of Different NN-based RadHAR Models Evaluated With Three Different Radar Settings alongside Their Model Size and Floating-point Operations (#\#FLOP.)
 Dataset DIAT [25]
@@
 (FMCW, Non-continuous)
-Dataset UoG2020 [27]
+Dataset UoG20 [27]
 (FMCW, Continuous)
@@
 #params
-#FLOP/Inf. a
+#\#FLOP a
 Accuracy
+Peak Mem
+Inf. Time
 #params
-#FLOP/Inf. a
+#\#FLOP. a
 Accuracy
+Peak Mem.
+Inf. Time
 #params
-#FLOP/Inf. a
+#\#FLOP a
 Accuracy
-ResNetb [35] 69.4 253.3 99.8$\pm$0.16 150.7 310.2 32.3$\pm$3.35 24.7 72.8 83.6$\pm$1.73
-Bi-LSTM [27] 23.3 5.6 93.9$\pm$0.68 134.4 6.0 79.4$\pm$1.23 15.5 2.1 85.3$\pm$2.01
-CNN-LSTM [13] 115.3 8.7 98.4$\pm$0.33 96.5 8.7 86.1$\pm$1.32 10.3 1.2 86.1$\pm$2.22
-CNN-Bi-GRU [14] 55.0 1090.1 98.5$\pm$0.41 142.8 1013.0 84.9$\pm$1.11 53.6 1082.0 85.4$\pm$2.19
-FML-ViTc [16] - - - $\sim$2700 169.0 $\sim$92.0 - - -
-LW-VITc [21] - - - 769 $\sim$2410 $\sim$92.0 - - -
-ActivityMambac [19] $\sim$8700.0 $\sim$1220.0 99.8 - - - - -
-RadMamba (Ours) 21.7 145.6 99.8$\pm$0.12 71.4 8.8 91.2$\pm$2.57 6.7 1.0 89.3$\pm$1.38
-RadMamba (Ours) 10.1 14.5 98.2$\pm$0.46 47.9 6.0 88.8$\pm$4.58 17.7 1.8 89.6$\pm$0.78
+Peak Mem.
+Inf. Time
+[…]
+[…]
+[…]
+[…]
+FML-ViTc [19] - - - - - $\sim$2700 169.0 $\sim$92.0 - - - - - - -
+LW-VITc [33] - - - - - 769 $\sim$2410 $\sim$92.0 - - - - - - -
+ActivityMambac [22] $\sim$8700.0 $\sim$1220.0 99.8 - - - - - - - - - - - -
+[…]
+[…]
   • a
-    ##\##FLOP/Inf. is the number of Floating point operations per inference.
+    #\#FLOP. is the number of floating-point operations per inference.
   • b
-    This ResNet only retains the structure of ResNet18 from the first layer to the end of the second residual block.
-The relative multiples of channels between convolution layers remain unchanged, but the size of the minimum channel can be adjusted.
+    This ResNet6 only retains the first 5 layers and the output linear layers of ResNet18.
   • c
     The $\sim$’ indicates that these results are approximated from the references [19, 33, 22].
-Figure 5: Comparison between ground truth and best prediction of RadMamba model with the configuration in Table I on (a) Dataset DIAT; (b) Dataset CI4R; (c) Dataset UoG2020.
+Figure 5: Example of comparisons between ground truth and best prediction of RadMamba model with the configuration in Table I on (a) Dataset DIAT; (b) Dataset CI4R; (c) Dataset UoG20.
+The experimental results detailed in Table V highlight the superior performance of our proposed RadMamba model across multiple radar-based human activity recognition datasets.
+Figure 5 (a) and (b) present confusion matrices comparing the ground truth labels with the best predictions of RadMamba, illustrating its classification accuracy and error distribution for each dataset.
+Figure 5 (c) shows the comparison between the classification results of RadMamba and CNN-LSTM over time.
+[…]
+For Dataset CI4R, employing FMCW radar in a non-continuous scenario with 11 activities, RadMamba attained an accuracy of 91.2% with a standard deviation of $\pm$2.57%, leveraging 71.4 thousand parameters.
+This performance outperforms previous models shown in Tab.
+V.
+Moreover, it remains competitive with the state-of-the-art of previous models while requiring substantially fewer computational resources (8.8 M #\#FLOP/Inf.
+vs.
+169 M to 2410 M #\#FLOP/Inf., showcasing its ability to excel in more complex activity recognition tasks with minimal computational overhead.
+Moreover, compared to the Bi-LSTM model, which has the lowest arithmetic intensity among the baselines, RadMamba improves accuracy by 9.4 percentage points while using 86.5 k fewer parameters and the same 6.0 M #FLOP/Inf.
+Figure 6: Accuracy vs. number of parameters (1) and number of #\#FLOP/Inf. (2) on: (a) Dataset DIAT (b) Dataset CI4R (c) Dataset UoG20
+On Dataset UoG20, which involves FMCW radar in a continuous setting with 6 activities, RadMamba recorded an accuracy of 89.3% with a standard deviation of $\pm$1.38%, utilizing 6.7 thousand parameters and 1.0M #\#FLOP/Inf.
+per inference.
+With an even smaller model size and a more complicated task than in Datasets DIAT and CI4R, RadMamba achieves the best accuracy among the models shown in Tab.
+V and highlights its effectiveness in continuous HAR scenarios with the smallest model size and #\#FLOP/Inf.
+per inference.
+[…]
+We also measured peak memory consumption to address concerns about efficiency.
+The results in Tab.
+V reveal relatively small differences between models, with peak CUDA allocated memory ranging from 9.7 MB to 22.8 MB.
+This limited variation, despite RadMamba’s significantly smaller parameter count, highlights a limitation of benchmarking on general-purpose, throughput-oriented embedded GPU architectures for edge inference when the batch size equals 1.
+On such platforms, the memory footprint is dominated by system-level overheads, such as CUDA context initialization, driver memory pools, and memory management, which mask the model’s intrinsic parameter efficiency.
+The true advantage of RadMamba’s parameter efficiency is its suitability for future hardware-software co-design.
+Its minimal footprint, ranging from 6.7k to 71.4k parameters across the three tested datasets, is intended to enable the entire model to be buffered within on-chip SRAM on a dedicated RadHAR accelerator.
+This approach would eliminate power-hungry off-chip DDR memory accesses, a primary source of power consumption in on-sensor systems.
+This fundamental power-saving benefit, along with the low arithmetic utilization for batch-size-1 inference on GPUs, cannot be captured by measurements on the Jetson platform.
+Therefore, we contend that the #\#params and #\#FLOPs metrics reported in Tab.
+V are more representative indicators of the RadMamba’s potential for enabling ultra-low-power, on-sensor inference, which is the primary goal of this work.
+VI-B Model Scaling and Efficiency Analysis
+The model size swept results, illustrated in Figure 6, compare accuracy against the number of parameters from 2k to 150k and #\#FLOP/Inf.
+across Datasets DIAT, CI4R, and UoG20 between previous models ResNet6, Bi-LSTM, CNN-Bi-GRU, and CNN-LSTM and the proposed RadMamba.
+For Dataset DIAT, as presented in Fig.
+6 (a1) and (a2), RadMamba achieves a peak accuracy of 99.8% with only 21.7 k parameters.
+However, RadMamba overfits with a model size larger than 21.7 k because the dataset is very simple and the accuracy is close to 100%.
+In contrast, ResNet6’s accuracy increases steadily with model size, reaching 99.9% with 208.1 k parameters, yet it requires nearly ten times the parameters of RadMamba for slightly lower performance.
 On Dataset CI4R, as shown in Fig.
-4 (b1) and (b2), RadMamba consistently outperforms previous models across a range of parameter configurations in efficiency, maintaining superior accuracy with fewer parameters and ##\##FLOP/Inf.
-Similarly, for Dataset UoG2020 shown in Fig.
-4 (c1) and (c2), RadMamba still records a higher accuracy for continuous activities scenarios with fewer ##\##FLOP/Inf., outperforming larger and more computationally intensive models in efficiency and effectiveness.
-These trends, as depicted in the figure, affirm RadMamba’s capability to achieve superior accuracy across diverse radar modalities while drastically reducing both parameter count and ##\##FLOP/Inf., positioning it as an optimal solution for lightweight, real-time human activity recognition.
-VII-B Comparison with Previous Works
-Figure 6: Averaged test loss across 10 random seeds vs. Epochs during training of models with configuration in Table I on (a) Dataset DIAT; (b) Dataset CI4R; (c) Dataset UoG2020.
-The experimental results detailed in Table IV highlight the superior performance of our proposed RadMamba model across multiple radar-based human activity recognition datasets.
-On Dataset DIAT, which utilizes CW radar in a non-continuous setting, RadMamba achieved an accuracy of 99.8% with a standard deviation of $\pm$0.12%, using only 21.7k parameters and requiring 145.6 million ##\##FLOP/Inf.
-per sample.
-This demonstrates its capability to match the state-of-the-art accuracy of previous models while being significantly more parameter-efficient compared to $\sim$8700k parameters and computationally lighter compared to $\sim$1220M ##\##FLOP per inference.
-The CNN-LSTM with 8.7 M ##\##FLOP per inference achieves better accuracy of 98.4% than 98.2% from RadMamba with 14.5 M ##\##FLOP per inference.
-However, take 45nm technology for example, the energy consumption for 8kB SRAM memory access of one 32-bit floating point parameter and one 32-bit floating point computing operation is 5.0 pJ and 4.6 pJ [36].
-With 11$\times$ larger model and around half ##\##FLOP per inference, it’s hard for CNN-LSTM to be more energy-efficient than RadMamba.
-TABLE V: Ablation Study with progressive techniques (denoted by Tech.) of New Blocks in RadMamba across Seed 0 to 5, Evaluated with Three Different Radar Settings alongside Their Number of Parameters
-Dataset DIAT Dataset CI4R Dataset UoG2020
+6 (b1) and (b2), RadMamba consistently outperforms previous models across a range of parameter configurations in efficiency, maintaining superior accuracy with fewer parameters and #\#FLOP/Inf.
+Similarly, for Dataset UoG20 shown in Fig.
+6 (c1) and (c2), RadMamba still records a higher accuracy for continuous activities scenarios with fewer #\#FLOP/Inf., outperforming larger and more computationally intensive models in efficiency and effectiveness.
+These trends, as depicted in the figure, affirm RadMamba’s capability to achieve superior accuracy across diverse radar modalities while drastically reducing both parameter count and #\#FLOP/Inf., positioning it as an optimal solution for lightweight, real-time human activity recognition.
+VI-C Ablation study of the Micro-doppler-oriented Structure
+TABLE VI: Ablation Study with progressive techniques (denoted by Tech.) of New Blocks in RadMamba across Seed 0 to 5, Evaluated with Three Different Radar Settings alongside Their Number of Parameters
+Dataset DIAT Dataset CI4R Dataset UoG20
 Row Tech. 1: Proj. Layer Tech. 2: Patch Size
@@
   • a
-    $(H_{1},W_{1})$ equals to (7, 7) for Dataset DIAT and B, (5, 7) for Dataset UoG2020 due to the different image size of these datasets.
+    $(H_{1},W_{1})$ equals to (7, 7) for Dataset DIAT and B, (5, 7) for Dataset UoG20 due to the different image size of these datasets.
   • b
@@
   • c
-    The kernel size for projection 1 and 2 is $(dim,3,3)$, and for projection 3 is $(dim,1,1)$.
-For Dataset CI4R, employing FMCW radar in a non-continuous scenario with 11 activities, RadMamba attained an accuracy of 91.2% with a standard deviation of $\pm$2.57%, leveraging 71.4 thousand parameters.
-This performance outperforms previous models shown in Tab.
-IV.
-Moreover, it remains competitive with the state-of-the-art of previous models while requiring substantially fewer computational resources (8.8 M ##\##FLOP/Inf.
-vs.
-169 M to 2410 M ##\##FLOP/Inf.), showcasing its ability to excel in more complex activity recognition tasks with minimal computational overhead.
-Moreover, compared to the Bi-LSTM model with lowest arithmetic intensity in table, RadMamba outperforms Bi-LSTM accuracy of 9.4% with 86.5 k less parameters and same ##\##FLOP per inference of 6.0 M.
-On Dataset UoG2020, which involves FMCW radar in a continuous setting with 6 activities, RadMamba recorded an accuracy of 89.3% with a standard deviation of $\pm$1.38%, utilizing 6.7 thousand parameters and 1.3M ##\##FLOP per inference.
-With an even smaller model size and more complicated task compared to Datasets DIAT and CI4R, RadMamba achieves the best accuracy in models shown in Tab.
-IV and highlights its effectiveness in continuous HAR scenarios with the smallest model size and ##\##FLOP per inference.
-Figure 5 (a) and (b) present confusion matrices comparing the ground truth labels with the best predictions of RadMamba, illustrating its classification accuracy and error distribution for each dataset.
-Figure 5 (c) shows the comparison between the classification results of RadMamba and CNN-LSTM over time.
-Complementing this, Figure 6 displays the averaged test loss across ten random seeds as a function of training epochs, providing insight into the model’s convergence behavior and stability during training on the same datasets.
-VII-C Ablation study of the Micro-doppler-oriented Structure
-Table V presents an ablation study evaluating the impact of new architectural blocks in the RadMamba model across three datasets with varying radar settings.
-The study systematically assesses the effects of different projection layers, patch sizes, and down-sampling scales on model accuracy and parameter efficiency, with results averaged over seeds 0 to 5.
-At the outset, the most important conclusion is that optimal performance emerges from the combined integration of all three proposed new blocks—projection layer, patch size, and down-sampling scale—rather than any single component in isolation.
+    The kernel size for projection 1 and 2 is $(dim,dim,3)$, and for projection 3 is $(dim,dim,1)$.
+Table VI presents an ablation study evaluating the impact of new architectural blocks in the RadMamba model across three datasets with varying radar settings.
+The study systematically assesses the effects of different projection layers, patch sizes, and downsampling scales on model accuracy and parameter efficiency, with results averaged over seeds 0 to 5.
+The main conclusion from this ablation study is that optimal performance emerges from the combined use of all three proposed new blocks, including the projection layer, the patch size, and the downsampling scale, rather than any single component in isolation.
 This is evident in the incremental improvements observed across the configurations.
-For instance, the baseline conventional ViM model (Row 1) achieves an accuracy of 93.5% on Dataset DIAT, 61.9% on Dataset CI4R, and 71.2% on Dataset UoG2020.
-In contrast, the proposed RadMamba (Row 27), which incorporates a 1D convolution layer, a patch size of ($H_{c-d}$, 1), and a down-sampling (2, 2), achieves a higher accuracy of 99.8% on Dataset DIAT.
-Similarly, on Dataset CI4R, Row 27 with a (2, 8) down-sampling scale yields significantly higher accuracy of 89.6%, and on Dataset UoG2020 with a (2, 32) scale, it reaches 88.6%.
-Intermediate configurations, such as Row 19 (only projection replacement), Row 7 (only patch size changing) and Row 2 (only down-sampling applied) show lower accuracies—underscoring that no single structure alone achieves the peak performance of the fully combined design in Row 27.
-To provide a more detailed analysis, we discuss the three blocks—projection layer, patch size, and down-sampling scale—separately, evaluating their individual contributions to RadMamba’s performance across the datasets.
-VII-C1 Projection Replacement
+For instance, the baseline conventional ViM model (Row 1) achieves an accuracy of 93.5% on Dataset DIAT, 61.9% on Dataset CI4R, and 71.2% on Dataset UoG20.
+In contrast, the proposed RadMamba (Row 27), which incorporates a 1D convolution layer, a patch size of ($H_{c-d}$, 1), and a downsampling (2, 2), achieves a higher accuracy of 99.8% on Dataset DIAT.
+Similarly, on Dataset CI4R, Row 27 with a (2, 8) downsampling scale yields significantly higher accuracy of 89.6%, and on Dataset UoG20 with a (2, 32) scale, it reaches 88.6%.
+Intermediate configurations, such as Row 19 (only projection replacement), Row 7 (only patch size changing), and Row 2 (only downsampling applied) show lower accuracies—underscoring that no single structure alone achieves the peak performance of the fully combined design in Row 27.
+To provide a more detailed analysis, we discuss the three blocks—projection layer, patch size, and downsampling scale—separately, evaluating their individual contributions to RadMamba’s performance across the datasets.
+VI-C1 Projection Replacement
 The adoption of a 1D convolutional layer as the projection mechanism consistently outperforms configurations with a single linear layer or three linear layers across nearly all settings.
-Comparing Rows 19-27 with Rows 10-18 and Rows 1-9, the conv1d projection replacement within the same patch size and down-sampling scale achieved average increment of mean accuracy by 0.48% over 1 linear layer, 0.40% over 3 linear layer on Dataset DIAT.
-On Dataset CI4R and UoG2020, the conv1d layer shows a substantial improvement of 2.99% and 3.64% over 1 linear layer, respectively.
-VII-C2 Patch Size
+Comparing Rows 19-27 with Rows 10-18 and Rows 1-9, the conv1d projection replacement within the same patch size and downsampling scale achieved an average increment of mean accuracy by 0.48% over 1 linear layer, 0.40% over 3 linear layers on Dataset DIAT.
+On datasets CI4R and UoG20, the conv1d layer shows a substantial improvement of 2.99% and 3.64% over 1 linear layer, respectively.
+VI-C2 Patch Size
 The ($H_{c-d}$, 1) configuration consistently delivers superior performance within each projection layer type.
-For the 1D convolutional layer, Row 25 with ($H_{c-d}$, 1) and without down-sampling achieves 3% gain on Dataset DIAT, 22.4% gain on Dataset CI4R, 9.2% gain on Dataset UoG2020 vs.
+For the 1D convolutional layer, Row 25 with ($H_{c-d}$, 1) and without downsampling achieves 3% gain on Dataset DIAT, 22.4% gain on Dataset CI4R, 9.2% gain on Dataset UoG20 vs.
 Row 19 with rectangular patch.
 This trend holds across other projection types, indicating that the ($H_{c-d}$, 1) patch size effectively retains the critical spatial features for radar data processing.
-VII-C3 Down-sampling
-The results suggest that down-sampling along the patch vector dimension generally outperforms down-sampling along the dimension orthogonal to patch vector and no down-sampling.
-For instance, Row 27’s (2, 32) time down-sampling achieves 88.6% $\pm$ 1.46, outperforming Row 26’s (8, 2) at 84.4% $\pm$ 2.05 and Row 25’s (1, 1) at 84.5% $\pm$ 0.77 on Dataset UoG2020.
+VI-C3 Downsampling
+The results suggest that downsampling along the patch vector dimension generally outperforms downsampling along the dimension orthogonal to the patch vector and no downsampling.
+For instance, Row 27’s (2, 32) time downsampling achieves 88.6% $\pm$ 1.46, outperforming Row 26’s (8, 2) at 84.4% $\pm$ 2.05 and Row 25’s (1, 1) at 84.5% $\pm$ 0.77 on Dataset UoG20.
 This pattern highlights the advantage of downsampling in enhancing the micro-Doppler features.
-The ablation study underscores the efficacy of the proposed RadMamba design, with the highest performance achieved by integrating a 1D convolutional projection layer, a ($H_{c-d}$, 1) patch size, and time-dimension down-sampling (Row 27).
-VIII Conclusion
-RadMamba sets a new benchmark in RadHAR, delivering top-tier accuracy with minimal parameters and computational cost across diverse datasets.
+The ablation study underscores the efficacy of the proposed RadMamba design, with the highest performance achieved by integrating a 1D convolutional projection layer, a ($H_{c-d}$, 1) patch size, and time-dimension downsampling (Row 27).
+VII Conclusion
+RadMamba sets a new benchmark in RadHAR (Code: https://github.com/lab-emi/AIRHAR), delivering top-tier accuracy with minimal parameters and computational cost across diverse datasets.
 Its radar micro-Doppler-oriented design addresses domain-specific challenges, making it ideal for real-time, resource-constrained applications.
-Overall, RadMamba consistently outperforms previous neural network models like ResNet, Bi-LSTM, CNN-LSTM, and CNN-Bi-GRU, and competes with or exceeds advanced architectures like Vision Transformers and ActivityMamba, all while maintaining a remarkably low parameter count (ranging from 6.7 k to 71.4 k) and reduced computational complexity (1.3 M to 145.6 M ##\##FLOP/Inf.).
+Overall, RadMamba consistently outperforms previous neural network models like ResNet6, Bi-LSTM, CNN-LSTM, and CNN-Bi-GRU, and competes with or exceeds advanced architectures like Vision Transformers and ActivityMamba, all while maintaining a remarkably low parameter count (ranging from 6.7 k to 71.4 k) and reduced computational complexity (1.3 M to 145.6 M #\#FLOP/Inf.).
 These results emphasize RadMamba’s capability as a lightweight and high-performance solution for radar-based HAR across diverse radar modalities and activity recognition tasks.
-The micro-Doppler-oriented design, integrating channel fusion, Doppler-aligned segmentation, and convolutional projections, addresses the unique challenges of radar data, delivering high accuracy across CW and FMCW datasets, both continuous and non-continuous.
-Ablation studies confirm the synergy of its components, with no single element achieving comparable performance alone.
+The micro-Doppler-oriented design, which integrates channel fusion, Doppler-aligned segmentation, and convolutional projections, addresses the unique challenges of radar data, delivering to deliver high accuracy across both CW and FMCW datasets, both including continuous and non-continuous data.
+Ablation studies confirm the synergy of its components, with no single element achieving comparable performance on its own.
 Acknowledgments
-This work was partially supported by the European Research Executive Agency (REA) under the Marie Skłodowska-Curie Actions (MSCA) Postdoctoral Fellowship program, Grant No.
+This work is funded by the European Research Executive Agency (REA) under the Marie Skłodowska-Curie Actions (MSCA) Postdoctoral Fellowship program, Grant No.
 101107534 (AIRHAR).